### v1

# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

Abstract.
In this paper, we construct a recursive subhomogeneous decomposition for the Cuntz–Pimsner algebras obtained from breaking the orbit of a minimal Hilbert $C(X)$-bimodule at a subset $Y\subset X$ with non-empty interior. This generalizes the known recursive subhomogeneous decomposition for orbit-breaking subalgebras of crossed products by minimal homeomorphisms.
Key words and phrases:
Minimal homeomorphisms, $\mathrm{C}^{*}$-correspondences, Hilbert bimodules, Cuntz–Pimsner algebras, recursive subhomogeneous algebras
2020 Mathematics Subject Classification:
46L35, 46L85, 46H25, 37B05
KRS is funded by RVO: 67985840. Part of this work was carried out when KRS was funded by GAČR project 20-17488Y and when JAJ was partially supported by NRF 2018R1D1A1B07041172.
1. Introduction
Recursive subhomogeneous (RSH) algebras are a particularly tractable class of unital subhomogeneous $\mathrm{C}^{*}$-algebras, which are given by iterated pullbacks of the form $C(X,M_{n})$, for $X$ a compact Hausdorff space. This class of $\mathrm{C}^{*}$-algebras is rich and includes homogeneous $\mathrm{C}^{*}$-algebras and their direct sums, section algebras of locally trivial continuous fields over compact spaces with matrix fibres, the dimension drop intervals and matrix algebras over them, and the subalgebras of crossed products by minimal homeomorphisms given by breaking the orbit of the homeomorphism at a closed subset with non-empty interior (see [28] and Section 5).
Introduced by Phillips in [32], the similarity of recursive subhomogeneous algebras to direct sums of homogeneous $\mathrm{C}^{*}$-algebras allowed for the extension of constructions required for the study of $K$-theory, stable, real, and decomposition rank, and simple inductive limits of interest to the Elliott classification program [32, 31, 42]. Indeed, when a simple unital $\mathrm{C}^{*}$-algebra $A$ can be modelled as an inductive limit of separable recursive subhomogeneous $\mathrm{C}^{*}$-algebra with slow dimension growth then it is classifiable [13]. In fact, inductive limits of RSH algebras are known to exhaust the Elliott invariant [12].
If a $\mathrm{C}^{*}$-algebra $A$ is -isomorphic to an RSH algebra $B$, then we say that $B$ is an RSH decomposition of $A$. In general, an RSH decomposition (if it exists) may not be unique. In this paper, we provide RSH decompositions for certain orbit-breaking subalgebras of Cuntz–Pimsner algebras of Hilbert bimodules over commutative $\mathrm{C}^{*}$-algebras.
If $A\cong C(X)$ for a compact metrizable space $X$, then any Hilbert bimodule over $A$ which is finitely generated projective as a right Hilbert $A$-module and is left full, is necessarily given by a right Hilbert module of sections of a line bundle over $X$ with left action given by the right action of $C(X)$ after composing with a homeomorphism $\alpha:X\to X$ [2, Proposition 3.7]. We say that the homeomorphism $\alpha$ is twisted by the line bundle $\mathscr{V}$ and denote such a module by $\Gamma(\mathscr{V},\alpha)$ and the corresponding Cuntz–Pimsner algebra as $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$. (Since $\Gamma(\mathscr{V},\alpha)$ is a Hilbert bimodule, one could equivalently construct the crossed product $C(X)\rtimes_{\Gamma(\mathscr{V},\alpha)}\mathbb{Z}$ in the sense of [1]. We use the Cuntz–Pimsner notation because it is slightly more concise.) When the vector bundle $\mathscr{V}$ is trivial, then $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is precisely the crossed product of $C(X)$ by the homeomorphism $\alpha$, $C(X)\rtimes_{\alpha}\mathbb{Z}$. Now let $Y\subset X$ be a closed non-empty subset. To this additional data we can associate the orbit-breaking Hilbert $C(X)$-module, $C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha)$, which is the Hilbert subbimodule of $\Gamma(\mathscr{V},\alpha)$ where we have restricted the left multiplication to those functions in $C(X)$ which vanish on the set $Y$. We call the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ an orbit-breaking algebra. When the vector bundle is trivial, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is isomorphic to an orbit-breaking subalgebra of the crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$.
From their introduction by Putnam in the case of Cantor minimal systems [34], orbit-breaking algebras have played a critical role in understanding properties and structures of crossed products such as $\mathcal{Z}$-stability and stable rank [26, 40, 41, 36, 35, 43, 25, 3]. They are also interesting $\mathrm{C}^{*}$-algebras in their own right, as sufficient conditions are known for when they are simple and $\mathcal{Z}$-stable. Here, $\mathcal{Z}$-stability refers to the property of tensorial absorption of the Jiang–Su algebra $\mathcal{Z}$—a simple, separable, unital, nuclear, projectionless, infinite-dimensional $\mathrm{C}^{*}$-algebra with unique tracial state and with the same $K$-theory as $\mathbb{Z}$ [21]. That is, a $\mathrm{C}^{*}$-algebra $A$ is $\mathcal{Z}$-stable if $A\otimes\mathcal{Z}\cong A$. As such, they provide dynamical models of $\mathrm{C}^{*}$-algebras that are covered by the classification theorem, which says that unital, simple, separable, nuclear $\mathrm{C}^{*}$-algebras which satisfy the Universal Coefficients Theorem (UCT) and are $\mathcal{Z}$-stable, are classified, up to -isomorphism, by $K$-theory and tracial data (see for example [9, 7, 14, 18, 17, 39, 8]). The UCT is, loosely speaking, a technical requirement for classification results that allows one to lift information from $KK$-theory to $K$-theory. All known nuclear $\mathrm{C}^{*}$-algebras satisfy the UCT, but it remains an open problem to determine whether nuclearity implies the UCT (see [6, 44] for a discussion).
We say that a $\mathrm{C}^{*}$-algebra unital, simple, separable, nuclear, $\mathcal{Z}$-stable $\mathrm{C}^{*}$-algebra which satisfies the UCT is classifiable. It was shown in [11] that orbit-breaking subalgebras of crossed products can realise a wide range of $K$-theory—much wider than the crossed products themselves—and hence a wide range of stably finite, classifiable $\mathrm{C}^{*}$-algebras. The introduction of a non-trivial line bundle has the potential to expand this range, possibly exhausting the range of the Elliott invariant in the stably finite case.
In [2], the authors introduced orbit breaking subalgebras of the Cuntz–Pimsner algebra of a minimal homeomorphism $\alpha$ over a compact metrizable space $X$, twisted by a vector bundle $\mathscr{V}$. When $\mathscr{V}$ is a line bundle and the subspace $Y\subset X$ at which the orbit broken is sufficiently small (for example, a finite set of points), it was proved that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is a centrally large subalgebra of $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ in the sense of Archey and Phillips [4]. This implies that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple and stably finite. Moreover, if the underlying space has finite dimension, it was shown that orbit breaking subalgebras are $\mathcal{Z}$-stable. To show this, one uses results about the Rokhlin dimension of a homeomorphism [38] on a finite-dimensional metric space to first show that the Cuntz–Pimsner algebra itself is $\mathcal{Z}$-stable. Again, appealing to the fact that it is centrally large, this means that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is also $\mathcal{Z}$-stable.
This implies, via the classification machinery, that $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is an inductive limit of RSH algebras. However, the result doesn’t say anything about the RSH building blocks in the inductive system. Even when $\mathscr{V}$ is trivial, and we are in the setting of crossed products and their orbit-breaking algebras, this result does not recover the approximately RSH structure given by Q. Lin in an unpublished preprint [27] (see [28] for an outline of the construction). It also gives no information on what happens when the subset $Y$ has non-empty interior (in this case the orbit-breaking algebra will not be simple).
The purpose of this paper is to exhibit an RSH decomposition for the Cuntz–Pimsner algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ when $Y\subset X$ is a closed subset with non-empty interior. This implies that when $Y$ is, for example, a single point, $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is approximately RSH.
Such a construction will allow us to investigate further properties of the $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles and their orbit-breaking subalgebras. In particular, a follow-up paper by the authors will show that if $(X,\alpha)$ is a minimal dynamical system with mean dimension zero (a dynamical generalization of covering dimension introduced by Gromov [19] and developed by Lindenstrauss and Weiss [29]), $\mathscr{V}$ is a line bundle, and $Y\subset X$ is sufficiently small, then both $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ are $\mathcal{Z}$-stable and hence classifiable. In subsequent work, further properties such as $K$-theory and stable rank will be studied.
The structure of the paper is as follows. Section 2 provides a short introduction to Hilbert bimodules and Cuntz–Pimsner algebras, and related preliminaries.
In Section 3 the tensor products of Hilbert $C(X)$-bimodules are described in terms of tensor products of line bundles and the composition of homeomorphisms. We make a considerable use of this description in this work.
In Section 5, we explain some properties of orbit-breaking algebras which are the Cuntz–Pimsner algebras of orbit-breaking bimodules. Then after briefly recalling the recursive subhomogeneous structure of orbit-breaking algebras in the case of crossed products by minimal dynamical homeomorphisms, we outline the construction of the RSH decomposition of orbit-breaking algebras.
Endomorphism bundles of vector bundles are the focus of Section 6. Vector bundles constructed from the line bundles in Section 3 are used in this framework to construct the building blocks of the recursive subhomogeneous.
Section 7 establishes an embedding of the orbit-breaking algebras into a direct of section algebras of endomorphism bundles.
The final section is devoted to the RSH decomposition for the range of the embedding obtained in Section 8. This implies that orbit-breaking algebras are recursive subhomogeneous algebras, which is the ultimate goal of this work.
Acknowledgments
The authors would like to thank Gi Hyun Park, Maria Grazia Viola, Magdalena Georgescu and Dawn Archey for helpful discussions.
2. Preliminaries
In this section we review the basics of Hilbert bimodules, with particular focus on Hilbert $C(X)$-bimodules. We then recall the construction of Cuntz–Pimsner algebras, as well as some useful facts that we will require in subsequent sections.
2.1. Hilbert bimodules
We are particularly interested in Hilbert $C(X)$-bimodules, but we begin with recalling definition and some results about Hilbert bimodules.
Definition 2.1.
Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras. A Hilbert $A$-$B$ bimodule $\mathcal{E}$ is a left $A$-module, right $B$-module equipped with
  • $\bullet$
    a right $B$-valued inner product $\langle\cdot,\cdot\rangle_{B}$ giving $\mathcal{E}$ a right Hilbert $B$-module structure, and
  • $\bullet$
    a left $A$-valued inner product ,${}_{A}\langle\cdot,\cdot\rangle$ giving $\mathcal{E}$ a left Hilbert $A$-module structure,
such that compatibility condition
$\xi_{1}\langle\xi_{2},\xi_{3}\rangle_{B}=\,_{A}\!\langle\xi_{1},\xi_{2}\rangle\xi_{3},$
is satisfied for every $\xi_{1},\xi_{2},\xi_{3}\in\mathcal{E}$. If $A=B$ we call $\mathcal{E}$ a Hilbert $A$-bimodule.
A Hilbert $A$-$B$-bimodule $\mathcal{E}$ is left, respectively right, full if ,$\overline{{}_{A}\langle\mathcal{E},\mathcal{E}\rangle}=A$, respectively $\overline{\langle\mathcal{E},\mathcal{E}\rangle_{B}}=B$. Here $\overline{\langle\mathcal{E},\mathcal{E}\rangle}$ denotes the closed linear span of $\langle\xi_{1},\xi_{2}\rangle$, $\xi_{1},\xi_{2}\in\mathcal{E}$.
Any Hilbert $A$-$B$-bimodule may also be viewed as a special case of an $A$-$B$ $\mathrm{C}^{*}$-correspondence, that is, a right Hilbert $B$-module equipped with a structure map $\varphi:A\to\mathcal{L}(\mathcal{E})$, which is a -homomorphism from $A$ into the $\mathrm{C}^{*}$-algebra of adjointable operators $\mathcal{L}(\mathcal{E})$. The structure map of a Hilbert $A$-$B$-bimodule is just given by left multiplication. Note that not every $A$-$B$ $\mathrm{C}^{*}$-correspondence can be given the structure of a Hilbert $A$-$B$-bimodule.
2.1.1. Tensor products of Hilbert bimodules
Let $A,B$ and $C$ be $\mathrm{C}^{*}$-algebras. Given a Hilbert $A$-$B$ bimodule $\mathcal{E}$ and a Hilbert $B$-$C$-bimodule $\mathcal{F}$, the $B$-balanced tensor product $\mathcal{E}\otimes_{B}\mathcal{F}$ is a Hilbert $A$-$C$-bimodule with left $A$-valued inner product satisfying
${}_{\mathcal{E}\otimes\mathcal{F}}\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle={}_{\mathcal{E}}\langle\xi_{1}{}_{\mathcal{F}}\langle\eta_{1},\eta_{2}\rangle,\xi_{2}\rangle,\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F},$
and right $C$-valued inner product satisfying
$\langle\xi_{1}\otimes\eta_{1},\xi_{2}\otimes\eta_{2}\rangle_{\mathcal{E}\otimes\mathcal{F}}=\langle\eta_{1},\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}}\eta_{2}\rangle_{\mathcal{F}},\quad\xi_{1},\xi_{2}\in\mathcal{E},\eta_{1},\eta_{2}\in\mathcal{F},$
If $\mathcal{E}$ is a Hilbert $A$-module then we denote
$\mathcal{E}^{\otimes n}:=\underbrace{\mathcal{E}\otimes_{A}\cdots\otimes_{A}\mathcal{E}}_{n\text{ times}}.$
2.1.2. Right Hilbert $C(X)$-modules
When $A=C(X)$, we can construct Hilbert $A$-modules from vector bundles. For a good introduction to the theory of bundles, see [20]. Let $X$ be a compact metric space and $\mathscr{V}=[T,p,X]$ a vector bundle of rank $n$. (Here $T$ denotes the total space of the bundle, $X$ the base space, and $p:T\to X$ the projection map.) We denote the set of continuous sections of $\mathscr{V}$ by $\Gamma(\mathscr{V})$. Then $\Gamma(\mathscr{V})$ has a natural right $C(X)$-module structure given by multiplication of a section by a continuous function. n fact, it is a right Hilbert $C(X)$-module: Let $\mathcal{A}=\{(U_{j},h_{j})\}_{j=1}^{N}$ be an atlas of $\mathscr{V}$ and let $\gamma_{1},\dots,\gamma_{N}$ a partition of unity subordinate to $U_{1},\dots,U_{N}$. Then the following $C(X)$-valued inner product
$\langle\xi,\eta\rangle_{\Gamma(\mathscr{V})}(x):=\sum_{j=1}^{N}\gamma_{j}(x)\langle h_{j}^{-1}(\xi(x)),h_{j}^{-1}(\eta(x))\rangle_{\mathbb{C}^{n}}$ (1)
makes $\Gamma(\mathscr{V})$ into a right Hilbert $C(X)$-module. Note that the inner product given in (1) does not depend on the choice of atlas. Moreover, with pairwise orthonormal vectors $s_{1},\dots,s_{n}$ of $\mathbb{C}^{n}$, the continuous sections given by
$\xi_{j,k}(x):=\begin{cases}h_{j}(x,s_{k}\gamma_{j}(x)^{1/2}),&x\in U_{j}\\ 0,&x\notin U_{j}.\end{cases}$
generate $\Gamma(\mathscr{V})$. Indeed, it is easily checked that
$\xi=\sum_{j=1}^{N}\sum_{k=1}^{n}\xi_{j,k}\langle\xi_{j,k},\xi\rangle_{\Gamma(\mathscr{V})}$
for every $\xi\in\Gamma(\mathscr{V})$. Any continuous function $f\in C_{0}(U_{j})$ allows us to define a section $\xi\in\Gamma(\mathscr{V})$ by setting
$\xi(x):=\left\{\begin{array}[]{ll}h_{j}(x,f(x))&\text{if}\quad x\in U_{j}\\ 0&\text{otherwise}.\end{array}\right.$
By abuse of notation, when we refer to such a section $\xi(x)$ we will often write $h_{j}(x,f(x))$. If $\mathscr{V}$ is a line bundle, then the sections
$\eta_{j}(x):=h_{j}(x,\gamma_{j}(x)^{1/2})$ (2)
generate $\Gamma(\mathscr{V})$.
By the Serre–Swan theorem [37, Theorem 2] every finitely-generated projective right $C(X)$-module is of the form $\Gamma(\mathscr{V})$ for a vector bundle over $X$. Furthermore, if $\mathcal{E}$ is a right Hilbert $C(X)$-module, there is a unitary isomorphism $U:\mathcal{E}\to\Gamma(\mathscr{V})$, where $\Gamma(\mathscr{V})$ is equipped with an inner product as defined in (1) with respect to any choice of atlas for $\mathscr{V}$.
2.1.3. Homeomorphisms twisted by line bundles
We recall from [2] that, when a Hilbert $C(X)$-module is left and right full, it always arises from a homeomorphism twisted by a line bundle over $X$.
Let $\alpha:X\to X$ be a homeomorphism and let $\mathscr{V}=[T,p,X]$ be a line bundle. We denote the pullback bundle by $\alpha^{*}\mathscr{V}$: the fibre $(\alpha^{*}\mathscr{V})_{x}$ of $\alpha^{*}\mathscr{V}$ at $x\in X$ is the fibre $\mathscr{V}_{\alpha(x)}$ of $\mathscr{V}$ at $\alpha(x)$, and $(x,\lambda)\mapsto h_{\alpha(U)}(\alpha(x),\lambda)$, $\lambda\in\mathbb{C}$, is a chart of $\alpha^{*}\mathscr{V}|_{U}$ whenever $(\alpha(U),h_{\alpha(U)})$ is a chart of $\mathscr{V}|_{\alpha(U)}$.
We define the Hilbert $C(X)$-bimodule $\Gamma(\mathscr{V},\alpha)$ as follows. As a right Hilbert $C(X)$-module, we have
$\Gamma(\mathscr{V},\alpha)_{C(X)}\cong\Gamma(\mathscr{V}),$
and we define the left action of $C(X)$ using $\alpha$ by setting
$f\xi:=\xi f\circ\alpha,\quad f\in C(X),\ \xi\in\Gamma(\mathscr{V},\alpha).$
The left inner product is then necessarily given by
${}_{C(X)}\langle\xi_{1},\xi_{2}\rangle=\langle\xi_{2},\xi_{1}\rangle_{C(X)}\circ\alpha^{-1}.$
Furthermore, as a left Hilbert $C(X)$-module, we have
${}_{C(X)}\Gamma(\mathscr{V},\alpha)\cong\Gamma((\alpha^{-1})^{*}\mathscr{V}),$
see [2, Lemma 3.6]. We refer to $\Gamma(\mathscr{V},\alpha)$ as the Hilbert $C(X)$-bimodule obtained by twisting the homeomorphism $\alpha$ by the line bundle $\mathscr{V}$. We refer to $\mathscr{V}$ as the twist.
Proposition 2.2 ([2, Propostion 3.7]).
Suppose that $\mathcal{E}$ is a non-zero Hilbert $C(X)$-bimodule which is finitely-generated and projective as a right Hilbert $C(X)$-module. Then there exist a compact subset $Y\subset X$, line bundle $\mathscr{V}$ over $X$ and homeomorphisms $\alpha:X\to Y$, $\beta:Y\to X$ such that
$\mathcal{E}_{C(X)}\cong\Gamma(\mathscr{V}),\quad_{C(X)}\mathcal{E}\cong\Gamma(\beta^{*}\mathscr{V}),$
and $f\xi=\xi f\circ\alpha$ for every $f\in C(X)$ and every $\xi\in\mathcal{E}$. Moreover, if $\mathcal{E}$ is left full, then we may take $X=Y$ and $\mathcal{E}\cong\Gamma(\mathscr{V},\alpha)$.
2.1.4. Orbit-breaking bimodules $\mathcal{E}_{Y}$
The main Hilbert bimodules of focus of this paper are the orbit-breaking subbimodules $\mathcal{E}_{Y}$ of $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$, which we define below.
Definition 2.3.
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a homeomorphism and $\mathscr{V}$ a line bundle over $X$. Let $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$. For a non-empty closed subset $Y\subset X$, the orbit-breaking bimodule is defined to be the subbimodule
$\mathcal{E}_{Y}:=C_{0}(X\setminus Y)\mathcal{E}$
that is obtained by restricting the left action on $\mathcal{E}$ to functions vanishing on $Y$. Note that $\mathcal{E}_{Y}$ is neither left nor right full as a Hilbert $C(X)$-bimodule.
2.2. Cuntz–Pimsner algebras
Cuntz–Pimsner algebras were originally introduced by Pimsner [33] and subsequently generalized by Katsura [22]. The class of Cuntz–Pimsner algebras includes all Cuntz–Krieger algebras, crossed products by $\mathbb{Z}$, crossed products by partial homeomorphisms and several other naturally occurring $\mathrm{C}^{*}$-algebras, see [22, Section 3].
Definition 2.4 ([23, Definition 2.1]).
Let $\mathcal{E}$ be a $\mathrm{C}^{*}$-correspondence over a $\mathrm{C}^{*}$-algebra $A$ with structure map $\varphi_{\mathcal{E}}:A\to\mathcal{L}(\mathcal{E})$. A representation $(\pi,\tau)$ of $\mathcal{E}$ on a $\mathrm{C}^{*}$-algebra $B$ consists of a -homomorphism $\pi:A\to B$ and a linear map $\tau:\mathcal{E}\to B$ satisfying
  1. (1)
    $\pi(\langle\xi,\eta\rangle_{\mathcal{E}})=\tau(\xi)^{*}\tau(\eta)$, for every $\xi,\eta\in\mathcal{E}$;
  2. (2)
    $\pi(a)\tau(\xi)=\tau(\varphi_{\mathcal{E}}(a)\xi)$, for every $\xi\in\mathcal{E}$, $a\in A$.
Note that (1) and the $\mathrm{C}^{*}$-identity imply $\tau(\xi)\pi(a)=\tau(\xi a)$ for every $\xi\in\mathcal{E}$ and $a\in A$.
Let $\psi_{\tau}:\mathcal{K}(\mathcal{E})\to B$ be the -homomorphism defined on rank one operators by
$\psi_{\tau}(\theta_{\xi,\eta})=\tau(\xi)\tau(\eta)^{*},\qquad\xi,\eta\in\mathcal{E}.$
We say that the representation $(\pi,\tau)$ is covariant [23, Definition 3.4] if in addition
  1. (3)
    $\pi(a)=\psi_{\tau}(\varphi_{\mathcal{E}}(a))$, for every $a\in J_{\mathcal{E}}:=\varphi_{\mathcal{E}}^{-1}(\mathcal{K}(\mathcal{E}))$.
In [23, Section 4] Kastura constructs a universal covariant representation $(\pi_{u},\tau_{u})$ satisfying the following: for any covariant representation $(\pi,\tau)$ of $\mathcal{E}$, there exists a surjective -homomorphism $\rho\colon\mathrm{C}^{*}(\pi_{u},\tau_{u})\rightarrow\mathrm{C}^{*}(\pi,\tau)$ such that $\pi=\rho\circ\pi_{u}$ and $\tau=\rho\circ\tau_{u}$, where ${\mathrm{C}}^{*}(\pi,\tau)$ is the $\mathrm{C}^{*}$-subalgebra of $B$ generated by $\pi(A)$ and $\tau(\mathcal{E})$.
Definition 2.5 ([23, Definition 3.5], see also [1, Definition 2.4]).
Let $A$ be a $\mathrm{C}^{*}$-algebra and let $\mathcal{E}$ be a $\mathrm{C}^{*}$-correspondence over $A$. The Cuntz–Pimsner algebra of $\mathcal{E}$ over $A$, denoted $\mathcal{O}_{A}(\mathcal{E})$, is the $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}(\pi_{u},\tau_{u})$ generated by the universal covariant representation of $\mathcal{E}$. We write $\mathcal{O}(\mathcal{E})$ if the $\mathrm{C}^{*}$-algebra $A$ is understood.
When $\mathcal{E}$ is a Hilbert $A$-bimodule, there is an equivalent construction due to Abadie, Eilers and Exel [1], which they call the crossed product of $A$ by the Hilbert $A$-bimodule $\mathcal{E}$. By [2, Proposition 3.7] the $C(X)$-correspondences of the form $\Gamma(\mathscr{V},\alpha)$ can be given the structure of a Hilbert bimodule only when $\mathscr{V}$ is a line bundle.
Examples 2.6.
Let $\alpha$ be a homeomorphism on an infinite compact metrizable space $X$, $\mathscr{V}=[T,p,X]$ a line bundle, and $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$.
  1. (1)
    If $\mathscr{V}$ is the trivial line bundle, then $\mathcal{O}(\mathcal{E})$ is isomorphic to the usual crossed product, $\mathcal{O}(\mathcal{E})\cong C(X)\rtimes_{\alpha}\mathbb{Z}$.
  2. (2)
    At the other extreme, if $\alpha=\mathrm{id}_{X}$ is the trivial homeomorphism, then $\mathcal{O}(\mathcal{E})$ is a commutative $\mathrm{C}^{*}$-algebra, and actually $\mathcal{O}(\mathcal{E})\cong C(P)$ where $P$ denotes the total space of the principal circle bundle $\mathcal{P}$ associated to the line bundle $\mathscr{V}$. (For further details see Section 7 in [2].)
  3. (3)
    Let $Y\subset X$ be a non-empty closed subset of $X$ and $\mathcal{E}_{Y}:=C_{0}(X\setminus Y)\mathcal{E}$ be the orbit-breaking bimodule. If $\mathscr{V}$ is trivial, then $\mathcal{O}(\mathcal{E}_{Y})\cong\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ is the orbit-breaking subalgebra introduced by Putnam in [34] and further studied in [28, 10, 11] and elsewhere.
A representation $(\pi,\tau)$ of $\mathcal{E}$ is said to admit a gauge action if, for each $z\in\mathbb{T}$, there exists a -homomorphism $\beta_{z}:\mathrm{C}^{*}(\pi,\tau)\to\mathrm{C}^{*}(\pi,\tau)$ such that $\beta_{z}(\pi(a))=\pi(a)$ and $\beta_{z}(\tau(\xi))=z\tau(\xi)$ for every $a\in A$ and every $\xi\in\mathcal{E}$ [23, Definition 5.6]. The gauge-invariant uniqueness theorem for Cuntz–Pimsner algebras is stated as follows.
Theorem 2.7 ([23, Theorem 6.4]).
For a covariant representation $(\pi,\tau)$ of a $\mathrm{C}^{*}$-correspondence $\mathcal{E}$ over a $\mathrm{C}^{*}$-algebra $A$, the surjective -homomorphism $\rho:\mathcal{O}(\mathcal{E})\to\mathrm{C}^{*}(\pi,\tau)$ is an isomorphism if and only if $(\pi,\tau)$ is injective (that is, $\pi$ is injective) and admits a gauge action.
The gauge action induces a topological grading
$\mathcal{O}(\mathcal{E})=\overline{\bigoplus_{n\in\mathbb{Z}}}$
where $E_{n}:=\{a\in\mathcal{O}_{A}(\mathcal{E})\mid\beta_{z}(a)=z^{n}a\}.$ Evidently each $E_{n}$ is itself a Hilbert $A$-bimodule. For $n>0$, we have $E_{n}\cong\mathcal{E}^{\otimes n}$, while for $n=0$, $E_{0}\cong A$.
To describe $E_{n}$ for $n<0$, let
^=$\widehat{\mathcal{E}}=\{\hat{\xi}\mid\xi\in\mathcal{E}\},$
where $\hat{\xi}:\mathcal{E}\to A$ is an $A$-bilinear map defined by $\hat{\xi}(\eta)=\langle\xi,\eta\rangle_{\mathcal{E}}$, $\xi,\eta\in\mathcal{E}.$ Then $\hat{\xi}\in\mathcal{E}^{\dagger}$, the dual Hilbert $A$-bimodule consisting of bounded $A$-linear maps $\varphi:\mathcal{E}\to A$, with actions given by
$(a\phi b)(\xi)=ab^{*}\phi(\xi),\qquad a,b\in A,\,\phi\in\mathcal{E}^{\dagger}.$
When $\mathcal{E}$ is finitely-generated projective as both a left and right $A$-module, we have that =^:=$\mathcal{E}^{\dagger}=\widehat{\mathcal{E}}:=\{\hat{\xi}\mid\xi\in\mathcal{E}\}$, see the discussion following Theorem 1.3 in [30]. We make $\widehat{\mathcal{E}}$ into a Hilbert $A$-bimodule by defining
${}_{\widehat{\mathcal{E}}}\langle\hat{\xi_{1}},\hat{\xi_{2}}\rangle:=\langle\xi_{1},\xi_{2}\rangle_{\mathcal{E}},\qquad\langle\hat{\xi_{1}},\hat{\xi_{2}}\rangle_{\widehat{\mathcal{E}}}:={}_{\mathcal{E}}\langle\xi_{1},\xi_{2}\rangle,$
for every $\xi_{1},\xi_{2}\in\mathcal{E}$. For $n<0$, we have $E_{n}\cong\widehat{\mathcal{E}}^{\otimes-n}$.
Note that if $\mathcal{E}^{\dagger}\cong\widehat{\mathcal{E}}$, then $\mathcal{E}$ is invertible in the sense that
^$\widehat{\mathcal{E}}\otimes_{A}\mathcal{E}\cong\mathcal{E}\otimes_{A}\widehat{\mathcal{E}}\cong A.$
3. Tensor products of Hilbert $C(X)$-bimodules $\Gamma(\mathscr{V},\alpha)$
As tensor products of Hilbert $C(X)$-bimodules play an important role in the construction of the relevant Cuntz–Pimsner algebras, we discuss the relation between the tensor products of Hilbert $C(X)$-bimodules and the Hilbert $C(X)$-bimodule constructed from the tensor product of the underlying line bundles and the composition of homeomorphisms.
For a line bundle $\mathscr{V}=[T,p,X]$ and a homeomorphism $\alpha:X\to X$, denote by
$\mathscr{V}^{(0)}:=X\times\mathbb{C}$
the trivial line bundle over $X$, and for each $n\geq 1$, define a line bundle $\mathscr{V}^{(n)}$ by
$\mathscr{V}^{(n)}$ $:=(\alpha^{n-1})^{*}\mathscr{V}\otimes(\alpha^{n-2})^{*}\mathscr{V}\otimes\cdots\otimes\alpha^{*}\mathcal{\mathscr{V}}\otimes\mathscr{V}.$ (3)
Whenever $\mathscr{V}|_{\alpha^{i}(U)}$ is trivial for all $i=0,1,\dots,n-1$ (via $h_{\alpha^{i}(U)}$), we have a chart map
$(x,\lambda)\mapsto\lambda h_{\alpha^{n-1}(U)}(\alpha^{n-1}(x),1)\otimes\cdots\otimes h_{U}(x,1)$
of the vector bundle $\mathscr{V}^{(n)}$ over $U$.
Note that the fibre $\mathscr{V}^{(n)}_{x}$ at $x\in X$ can be written as
$\mathscr{V}_{x}^{(n)}=\mathscr{V}_{\alpha^{m}(x)}^{(n-m)}\otimes\mathscr{V}_{x}^{(m)}$ (4)
whenever $0\leq m\leq n$. In fact,
$\mathscr{V}_{x}^{(n)}$ $=\mathscr{V}_{\alpha^{n-1}(x)}\otimes\mathscr{V}_{\alpha^{n-2}(x)}\otimes\cdots\otimes\mathscr{V}_{\alpha^{m}(x)}\otimes\mathscr{V}_{\alpha^{m-1}(x)}\otimes\cdots\otimes\mathscr{V}_{\alpha(x)}\otimes\mathscr{V}_{x}$
$=\mathscr{V}_{\alpha^{n-1-m}(\alpha^{m}(x))}\otimes\cdots\otimes\mathscr{V}_{\alpha^{m}(x)}\otimes\mathscr{V}_{\alpha^{m-1}(x)}\otimes\cdots\otimes\mathscr{V}_{\alpha(x)}\otimes\mathscr{V}_{x}$
$=\mathscr{V}_{\alpha^{m}(x)}^{(n-m)}\otimes\mathscr{V}_{x}^{(m)}.$
The following result is surely well known when the homeomorphism $\alpha$ is trivial and should be in the literature, however we could not find a reference.
Proposition 3.1.
Let $\alpha:X\to X$ be a homeomorphism on a compact metric space $X$ and $\mathscr{V}$ be a line bundle over $X$. Then there is a Hilbert $C(X)$-bimodule isomorphism
$\psi:\Gamma(\mathscr{V},\alpha)^{\otimes n}\to\Gamma(\mathscr{V}^{(n)},\alpha^{n})$
such that for $\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n}\in\Gamma(\mathscr{V},\alpha)^{\otimes n}$, we have
$\psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n})(x)=\,\xi_{1}(\alpha^{n-1}(x))\otimes\xi_{2}(\alpha^{n-2}(x))\otimes\cdots\otimes\xi_{n}(x).$
If $n=0$, then by $\psi$ we mean the identity map on $C(X)$.
Proof.
If $\xi_{i}\in\Gamma(\mathscr{V})$, $i=1,\dots,n$, then clearly the map
$x\mapsto\xi_{1}(\alpha^{n-1}(x))\otimes\cdots\otimes\xi_{n-1}(\alpha(x))\otimes\xi_{n}(x),\ x\in X,$
is a continuous section in $\Gamma(\mathscr{V}^{(n)})$, and thus we have a multilinear map sending $(\xi_{1},\dots,\xi_{n})\in\Gamma(\mathscr{V})\times\cdots\times\Gamma(\mathscr{V})$ to this section in $\Gamma(\mathscr{V}^{(n)})$. Then there is a linear map $\psi:\Gamma(\mathscr{V})^{\otimes n}\to\Gamma(\mathscr{V}^{(n)})$ satisfying the identity in the proposition. For $f\in C(X)$,
$\psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n}f)(x)=(\psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n})f)(x)$
is rather obvious, and also we have
$\ \psi(f\cdot\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n})(x)$
$=$ $\ \psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n}f\circ\alpha^{n})(x)$
$=$ $\ \xi_{1}(\alpha^{n-1}(x))\otimes\cdots\otimes\xi_{n-1}(\alpha(x))\otimes\xi_{n}(x)f\circ\alpha^{n}(x)$
$=$ $\ f\cdot\psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n})(x).$
Thus $\psi$ is a $C(X)$-bimodule map from $\Gamma(\mathscr{V},\alpha)^{\otimes n}$ to $\Gamma(\mathscr{V}^{(n)},\alpha^{n})$.
To show that $\psi$ preserves the inner product, with $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$, we calculate
$\,\langle\xi_{1}\otimes\cdots\otimes\xi_{n},\eta_{1}\otimes\cdots\otimes\eta_{n}\rangle_{\mathcal{E}^{\otimes n}}(x)$
$=$ $\langle\xi_{2}\otimes\cdots\otimes\xi_{n},\,\langle\xi_{1},\eta_{1}\rangle_{\mathcal{E}}\eta_{2}\otimes\cdots\otimes\eta_{n}\rangle_{\mathcal{E}^{\otimes n-1}}(x)$
$=$ $\,\langle\xi_{2}\otimes\cdots\otimes\xi_{n},\eta_{2}\otimes\cdots\otimes\eta_{n}\rangle_{\mathcal{E}^{\otimes n-1}}\langle\xi_{1},\eta_{1}\rangle_{\mathcal{E}}\circ\alpha^{n-1}(x)$
[…]
$=$ $\,\langle\xi_{n},\eta_{n}\rangle_{\mathcal{E}}(x)\langle\xi_{n-1},\eta_{n-1}\rangle_{\mathcal{E}}\circ\alpha(x)\cdots\langle\xi_{1},\eta_{1}\rangle\circ\alpha^{n-1}(x),$
and
$\,\langle\psi(\xi_{1}\otimes\xi_{2}\otimes\cdots\otimes\xi_{n}),\psi(\eta_{1}\otimes\eta_{2}\otimes\cdots\otimes\eta_{n})\rangle_{\Gamma(\mathscr{V}^{(n)})}(x)$
$=$ $\,\langle\xi_{1}(\alpha^{n-1}(x))\otimes\cdots\otimes\xi_{n}(x),\eta_{1}(\alpha^{n-1}(x))\otimes\cdots\otimes\eta_{n}(x)\rangle_{\mathscr{V}^{(n)}_{x}}$
$=$ $\,\langle\xi_{1}(\alpha^{n-1}(x)),\eta_{1}(\alpha^{n-1}(x))\rangle_{\mathscr{V}_{\alpha^{n-1}(x)}}\cdots\langle\xi_{n}(x),\eta_{n}(x)\rangle_{\mathscr{V}_{x}},$
and compare them. We see that these two match up to the order of multiplication since
$\langle\xi_{i},\eta_{i}\rangle_{\mathcal{E}}(\alpha^{n-i}(x))=\langle\xi_{i}(\alpha^{n-i}(x)),\eta_{i}(\alpha^{n-i}(x))\rangle_{\mathscr{V}_{\alpha^{n-i}(x)}}.$
This inner product preserving property implies that $\psi$ is injective.
Let $\eta_{j}$ be the generators of $\mathcal{E}=\Gamma(\mathscr{V})$ as in (2). From the definition of $\mathscr{V}^{(n)}$, one can show that the sections $\psi(\eta_{i_{1}}\otimes\eta_{i_{2}}\otimes\cdots\otimes\eta_{i_{n}})$ generate $\Gamma(\mathscr{V}^{(n)},\alpha^{n})$. Then for $\xi\in\Gamma(\mathscr{V},\alpha)^{\otimes n}$, the element
$\mu$ $:=\sum_{i_{1},i_{2},\ldots,i_{n}}\eta_{i_{1}}\otimes\eta_{i_{2}}\otimes\cdots\otimes\eta_{i_{n}}\langle\psi(\eta_{i_{1}}\otimes\eta_{i_{2}}\otimes\cdots\otimes\eta_{i_{n}}),\xi\rangle_{\Gamma(\mathscr{V}^{(n)})}$
in $\mathcal{E}^{\otimes n}$ satisfies $\psi(\mu)=\xi$. Thus $\psi$ is surjective. ∎
We end this section with the following proposition about orbit-breaking bimodules which will be used in the following sections.
Proposition 3.2 (see [2, Lemma 6.2]).
Let $\alpha:X\to X$ be a homeomorphism and $\mathscr{V}$ be a line bundle on $X$. Let $\psi:\mathcal{E}^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$ be the isomorphism given by Proposition 3.1. Let $Y\subset X$ be a non-empty closed subset and denote by $\mathcal{E}_{Y}$ the associated orbit-breaking bimodule. Then for any element $\xi\in\mathcal{E}^{\otimes m}$, we have that $\xi\in\mathcal{E}_{Y}^{\otimes m}:=(\mathcal{E}_{Y})^{\otimes m}$ if and only if $\psi(\xi)$ vanishes on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. If $m=0$, then this set is empty and by $\psi$ we mean the identity map of $C(X)$.
Proof.
If $\xi=f_{1}\eta_{1}\otimes\cdots\otimes f_{m}\eta_{m}$ for some $\eta_{i}\in\mathcal{E}$, $f_{i}\in C_{0}(X\setminus Y)$, $i=1,2,\ldots,m$, then
$\xi$ $=f_{1}\eta_{1}\otimes\cdots\otimes f_{m}\eta_{m}$
$=\eta_{1}\otimes\cdots\otimes\eta_{m}(f_{1}\circ\alpha^{m})\cdots(f_{m-1}\circ\alpha^{2})(f_{m}\circ\alpha),$
thus $\psi(\xi)(x)=0$ whenever $x\in\alpha^{-1}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since such $\xi$ generate $\mathcal{E}_{Y}^{\otimes m}$, the only if part follows.
Conversely, suppose $\xi\in\mathcal{E}^{\otimes m}$ satisfies $\psi(\xi)(x)=0$ for every $x\in\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Let $\{h_{j}:U_{j}\times\mathbb{C}\to\mathscr{V}|_{U_{j}}\}_{j=1,\ldots,n}$ be an atlas for $\mathscr{V}$ and $\gamma_{1},\ldots,\gamma_{n}$ a partition of unity subordinate to $U_{1},\ldots,U_{n}$. Then $\eta_{j}$, $1\leq j\leq n$, given by $\eta_{j}(x)=h_{j}(x,{\gamma_{j}}^{1/2}(x))$, generates $\mathcal{E}$ as a right Hilbert $C(X)$-module, hence $\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}$, $1\leq i_{1},\dots,i_{m}\leq n$ generates $\mathcal{E}^{\otimes m}$. Then the functions
$f_{i_{1},i_{2},\ldots,i_{m}}(x):=\langle\psi(\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}),\psi(\xi)\rangle_{\Gamma(\mathscr{V}^{(m)})}(x)=\langle\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}},\xi\rangle_{\mathcal{E}^{\otimes m}}(x)$
vanishes on $\alpha^{-1}(Y)\cup\alpha^{-2}(Y)\cup\cdots\cup\alpha^{-m}(Y)$. Since the functions $f_{i_{1},i_{2},\ldots,i_{m}}$ are linear combinations of positive ones, assuming that they are positive in the following computation, we have
$\xi$ $=\sum_{1\leq i_{1},\dots,i_{m}\leq n}\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}\langle\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}},\xi\rangle_{\mathcal{E}^{\otimes m}}$
$=\sum_{1\leq i_{1},\ldots,i_{m}\leq n}\eta_{i_{1}}\otimes\cdots\otimes\eta_{i_{m}}f_{i_{1},i_{2},\ldots,i_{m}}$
$=\sum_{1\leq i_{1},\ldots,i_{m}\leq n}\big{[}(f_{i_{1},i_{2},\ldots,i_{m}}\circ\alpha^{-m})^{1/m}\eta_{i_{1}}\otimes(f_{i_{1},i_{2},\ldots,i_{m}}\circ\alpha^{-m+1})^{1/m}\eta_{i_{2}}\otimes\cdots$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\cdots\otimes(f_{i_{1},i_{2},\ldots,i_{m}}\circ\alpha^{-1})^{1/m}\eta_{i_{m}}\big{]}.$
Thus $\xi\in\mathcal{E}_{Y}^{\otimes m}$. ∎
4. Orbit-breaking algebras
Generalizing Example 2.6.(3), we define orbit-breaking algebras via orbit-breaking bimodules. The definition of orbit-breaking algebras in this context first appeared in [2].
Definition 4.1.
Let $\alpha:X\to X$ be a homeomorphism on an infinite compact metric space $X$ and $\mathscr{V}$ be a line bundle over $X$. Let $\mathcal{E}=\Gamma(\mathscr{V},\alpha)$. Suppose that $Y\subseteq X$ is a non-empty closed subset of $X$. The orbit-breaking algebra of $\mathcal{O}(\mathcal{E})$ at $Y$ is the Cuntz–Pimsner algebra $\mathcal{O}(\mathcal{E}_{Y})$ of the orbit breaking submodule $\mathcal{E}_{Y}=C_{0}(X\setminus Y)\mathcal{E}$.
Note that the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ can be identified with the $\mathrm{C}^{*}$-subalgebra $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)\mathcal{E})$ of $\mathcal{O}(\mathcal{E})$.
Let $A$ be a $\mathrm{C}^{*}$-algebra and $\mathcal{E}$ a $\mathrm{C}^{*}$-correspondence over $A$ with the structure map $\varphi_{\mathcal{E}}$. We say that $\mathcal{E}$ is minimal if, for any non-zero ideal $J\subset A$, if $\langle\mathcal{E},\varphi_{E}(J)\mathcal{E}\rangle_{\mathcal{E}}\subset J$, then $J=A$. This definition is a generalization of the corresponding definition for a single homeomorphism $\alpha:X\to X$ on a compact metric space: We say that $\alpha$ is minimal if, for every non-empty closed subset $Y\subset X$, if $\alpha(Y)\subset Y$, then $Y=X$.
Theorem 4.2.
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a homeomorphism, and $\mathscr{V}$ a line bundle over $X$. Let $Y\subset X$ be a non-empty closed subset. Then
  1. (1)
    $\Gamma(\mathscr{V},\alpha)$ is minimal if and only if $\alpha$ is minimal if and only if the crossed product $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is simple.
  2. (2)
    If $\alpha$ ia minimal, then $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)=\emptyset$ for every $n\in\mathbb{Z}\setminus\{0\}$.
  3. (3)
    Both $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ are nuclear and satisfy the UCT.
Proof.
That $\Gamma(\mathscr{V},\alpha)$ is minimal if and only if $\alpha$ is minimal if and only if the crossed product $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ is simple follows from [2, Proposition 3.9, Corollary 3.11]. Suppose that $\alpha$ is minimal. That $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is simple if and only if $Y\cap\alpha^{n}(Y)\neq\emptyset$ for $n\neq 0$ follows from [24]. (Note that this is a slight improvement of the result stated in [2]). Applying [23, Theorem 7.3, Proposition 8.8] to Cuntz–Pimsner algebras $\mathcal{O}(\Gamma(\mathscr{V},\alpha))$ and $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ yields (3). ∎
5. Recursive subhomogeneous algebras
Let $X$ be an infinite compact metric space, $\alpha:X\to X$ a minimal homeomorphism, $\mathscr{V}$ a line bundle over $X$ and $Y\subset X$ a closed subset with non-empty interior. As the construction of a recursive subhomogeneous decomposition of an orbit-breaking algebra $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$ is quite technical, this section recalls the definition of a recursive subhomogeneous algebras as well as the recursive subhomogeneous algebra structure of orbit-breaking algebras for crossed products by minimal homeomorphisms. We then describe the conceptual steps in obtaining the recursive subhomogeneous of $\mathcal{O}(C_{0}(X\setminus Y)\Gamma(\mathscr{V},\alpha))$, which highlights the similarities and differences to the construction in the crossed product case.
Let $A$, $B$ and $C$ be $\mathrm{C}^{*}$-algebras and $\varphi:A\to C$ and $\rho:B\to C$ be homomorphisms. The pullback $A\oplus_{C}B$ is defined to be the algebra of all $(a,b)\in A\oplus B$ such that $\varphi(a)=\rho(b)$.
Definition 5.1 ([32]).
A recursive subhomogeneous algebra is a $\mathrm{C}^{*}$-algebra given by the following recursive definition.
  1. (1)
    If $X$ is a compact Hausdorff space and $n\geq 1$, then $C(X,M_{n})$ is a recursive subhomogeneous algebra.
  2. (2)
    If $A$ is a recursive subhomogeneous algebra, $X$ is a compact Hausdorff space, $X^{(0)}\subseteq X$ is closed, $\varphi:A\to C(X^{(0)},M_{n})$ is any unital homomorphism, and $\rho:C(X,M_{n})\to C(X^{(0)},M_{n})$ is the restriction homomorphism, then the pullback $A\oplus_{C(X^{(0)},M_{n})}C(X,M_{n})$ is a recursive subhomogenous algebra.
A recursive subhomogeneous decomposition for a given $\mathrm{C}^{*}$-algebra is in general not unique. Observe that any recursive subhomogeneous $\mathrm{C}^{*}$-algebra can be written in the form
$R=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{l}^{(0)}}C_{l},$ (5)
where $C_{j}=C(X_{j},M_{n(j)})$ for compact Hausdorff spaces $X_{j}$, $n(j)\in\mathbb{Z}_{>0}$ and $C^{(0)}_{j}=C(X_{j}^{(0)},M_{n(j)})$ for (possibly empty) compact subsets $X^{(0)}_{j}\subset X_{j}$, and where the maps $C_{j}\to C_{j}^{(0)}$ are the restriction maps. The expression (5) is called a recursive subhomogeneous decomposition of $R$.
Definition 5.2.
Let
$R=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{l}^{(0)}}C_{l}$
be a recursive subhomogeneous decomposition. Then
  1. (1)
    $l$ is called the length of $R$,
  2. (2)
    the recursive subhomogeneous algebra
    $R^{(k)}=\left[\cdots\left[\left[C_{0}\oplus_{C^{(0)}_{1}}C_{1}\right]\oplus_{C_{2}^{(0)}}\right]\cdots\right]\oplus_{C_{j}^{(0)}}C_{j},$
    is called the $k$th stage algebra,
  3. (3)
    $X_{0},\dots,X_{l}$ are called the base spaces of the decomposition and their disjoint union $\bigsqcup_{j=1}^{l}X_{j}$ is called the total space,
  4. (4)
    $n(0),\dots,n(j)$ are called the matrix sizes.
Let $\alpha$ be a minimal homeomorphism on a compact metrizable space $X$ and $Y$ be a closed subset of $X$ with non-empty interior as before. In the following, we will denote the orbit-breaking algebra $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)$ by $A_{Y}$. For $x\in Y$, the first return time $r_{Y}(x)$ of $x$ to $Y$ is the smallest integer $n\geq 1$ such that $\alpha^{n}(x)\in Y$. Then it can be checked that $\sup_{x\in Y}r_{Y}(x)<\infty$. Let $r_{1}<\dots<r_{K}$ be the distinct values of the first return times to Y and set
$Y_{k}=\{x\in Y\mid r_{Y}(x)=r_{k}\},$ (6)
for all $1\leq k\leq K$. If $Y$ is understood, we use $r(x)$ in place of $r_{Y}(x)$. Note that $\cup_{i=1}^{k}Y_{i}$ is always closed for every $k=1,\dots,K$, and
$Y=\bigsqcup_{k=1}^{K}Y_{k},\quad\mbox{and}\quad X=\bigsqcup_{k=1}^{K}\bigsqcup_{i=0}^{r_{k}-1}\alpha^{i}(Y_{k}).$ (7)
The sequences
$Y_{k},\alpha(Y_{k}),\cdots,\alpha^{r_{k}-1}(Y_{k})$
are called Rokhlin towers. The sets $Y_{k}$ are the bases of towers and $r_{k}$ are their heights.
If $X$ is a Cantor set, then each $Y_{k}$ is clopen and compact. Set $p_{k}=\chi_{Y_{k}}$, then each $p_{k}$ commutes with elements in $A_{Y}$ and we have $A_{Y}=\bigoplus_{k=1}^{K}p_{k}A_{Y}p_{k}$. Then $\pi_{k}\colon p_{k}A_{Y}p_{k}\to C(Y_{k},M_{r_{k}})$ given by
$\pi_{k}(f)=\mathrm{diag}(f|_{Y_{k}},\dots,f\circ\alpha^{r_{k}-1}|_{Y_{k}})$
for $f\in C(X)$ and
$\pi_{k}(\chi_{\alpha^{j}(Y_{k})}u\chi_{\alpha^{j-1}(Y_{k})})=e_{j,j-1}\otimes 1.$
Then each $\pi_{k}$ is an isomorphism, specifically, $A_{Y}\cong\bigoplus_{k=1}^{K}C(Y_{k},M_{r_{k}})$.
If $X$ is not a Cantor set, then $Y_{k}$’s are not necessarily closed and boundaries $\alpha^{j}(\overline{Y_{k}})$ can overlap which lead to a recursive subhomogeneous decomposition for orbit breaking subalgebras instead of direct limit of homogeneous algebras.
Set
$B_{Y}=\bigoplus_{k=1}^{K}C(\overline{Y_{k}},M_{r_{k}}).$
There exists an embedding
$\pi:A_{Y}\to B_{Y}.$
Let $\pi_{k}:A_{Y}\to C(\overline{Y_{k}},M_{r_{k}})$ be the composition of $\pi$ with the projection into the $k$-summand of $B_{Y}$. Set $\pi^{(k)}=\oplus_{i=1}^{k}\pi_{k}$ and $A_{k}=\pi^{(k)}(A_{Y})$. We say that an element $b=(b_{1},\dots,b_{K})$ has a boundary decomposition property if it satisfies the following: for any $x\in\overline{Y_{k}}\setminus Y_{k}$, $k=1,\dots,K$ with
$x\in\overline{Y_{k}}\cap Y_{t_{1}}\cap\alpha^{-{r_{t_{1}}}}(Y_{t_{2}})\cap\cdots\cap\alpha^{-(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m-1}})}(Y_{t_{m}})$
and $r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m}}=r_{k}$, then $b_{k}$ is given by the block diagonal matrix
$b_{k}(x)=\mathrm{diag}(b_{t_{1}}(x),b_{t_{2}}(\alpha^{r_{t_{1}}}(x)),\dots,b_{t_{s}}(\alpha^{r_{t_{1}}+\dots+r_{t{s-1}}}(x))).$
Then $b=(b_{1},\dots,b_{K})\in B_{Y}$ is in $\pi(A_{Y})$ if and only if $b$ has the boundary decomposition property. This result enables us to explicitly describe the recursive subhomogeneous construction of orbit-breaking subalgebra $A_{Y}$. Fix $1\leq k\leq K$ to define $\varphi_{k}:A_{k-1}\to C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})$, take arbitrary $(b_{1},\dots,b_{k-1})\in A_{k-1}$. By the definition there is $b=(b_{1},\dots,b_{k-1},b_{k},\dots,b_{K})\in\pi(A_{Y})$. Since $b$ satisfies the boundary decomposition property, $b_{k}|_{\overline{Y_{k}}\setminus Y_{k}}$ is only depend on $(b_{1},\dots,b_{k-1})$. Now, define $\varphi_{k}(b_{1},\dots,b_{k-1})(x)=b_{k}(x)$ and let $\rho_{k}:C(\overline{Y_{k}},M_{r_{k}})\to C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})$ denote the restriction map. Then
$A_{k}=A_{k-1}\oplus_{C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})}C(\overline{Y_{k}},M_{r_{k}}).$
It follows that $A_{Y}$ has a recursive subhomogeneous decomposition given by
$[\dots[C(Y_{1},M_{r_{1}})\oplus_{C(\overline{Y_{2}}\setminus Y_{2},M_{r_{2}})}C(\overline{Y}_{2},M_{r_{2}})]\dots]\oplus_{C(\overline{Y_{k}}\setminus Y_{k},M_{r_{k}})}C(\overline{Y_{k}},M_{r_{k}})]$ (8)
Now let $\mathscr{V}$ be a line bundle over $X$ and set $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$. We describe the conceptual steps in obtaining a recursive subhomogeneous decomposition for the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$.
Step 1. In Section 6, we begin with introducing the building blocks in the recursive subhomogeneous decomposition the of orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ Set $\mathscr{D}^{(r_{k})}=\mathscr{V}^{(0)}\oplus\dots\oplus\mathscr{V}^{(r_{k}-1)}$ where $\mathscr{V}^{(i)}$ is the line bundle defined in (3), and let $\mathscr{M}_{k}:=End(\mathscr{D}^{(r_{k})})|_{\overline{Y}_{k}}$ be the endomorphism bundle (see Section 6). We show in Theorem 8.7 that $C(\overline{Y}_{k},M_{r_{k}})$ in (8) will be replaced by $\Gamma(\mathscr{M}_{k})$ when the twist is not trivial.
Step 2. In Section 7, we construct an embedding from the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. For this, we introduce a covariant representation $(\pi_{k},\tau_{k})$ of the Hilbert $C(X)$-bimodule $\mathcal{E}_{Y}$ to $\Gamma(\mathscr{M}_{k})$ which induces a *-homomorphism, again denoted by $\pi_{k}$, from the orbit-breaking algebra to $\Gamma(\mathscr{M}_{k})$, for all $1\leq k\leq K$. In Proposition 7.4 we show that $\pi=\oplus_{k=1}^{K}\pi_{k}$ is injective and its image admits a gauge action, which, by the gauge invariant uniqueness theorem, tells us that $\pi$ is a -isomorphism from $\mathcal{O}(\mathcal{E}_{Y})$ onto its image.
Step 3. To describe the pullbacks in the RSH structure, we need to understand the image of the *-representation $\pi$. We generalize the boundary decomposition for elements in $\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$ in Definition 8.1. Then by theorem 8.6 the range of $\pi$ is precisely the set of all elements with boundary decomposition property.
Step 4. Finally we obtain the main result of the paper. The boundary decomposition property of elements in the image of $\pi$ enables us to define *-homomorphisms $\varphi_{k}:(\oplus_{i=1}^{k-1}\pi_{i})(\mathcal{O}(\Gamma(\mathcal{E}_{Y}))\to\Gamma(\mathscr{M}|_{\overline{Y}_{k}\setminus Y_{k}})$. Then by Theorem 8.7, the pullback $(\oplus_{i=1}^{k-1}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))\oplus_{\Gamma(\mathscr{M}_{k}|_{\overline{Y}_{k}\setminus Y_{k}})}\Gamma(\mathscr{M}_{k})$ via $\varphi_{k}$ and the restriction map is equal to $(\oplus_{i=1}^{k}\pi_{i})(\mathcal{O}(\mathcal{E}_{Y}))$. That is,
$\mathcal{O}(\mathcal{E}_{Y})\cong\ [\cdots[\Gamma(\mathscr{M}_{1})\oplus_{\Gamma(\mathscr{M}_{2}|_{\overline{Y}_{2}\setminus Y_{2}})}\Gamma(\mathscr{M}_{2})]\oplus\cdots]\oplus_{\Gamma(\mathscr{M}_{K}|_{\overline{Y}_{k}\setminus Y_{K}})}\Gamma(\mathscr{M}_{K}).$
Every $\Gamma(\mathscr{M}_{k})$ is itself an RSH algebra [32, Proposition 1.7], which gives us the RSH decomposition of the orbit-breaking subalgebra.
6. Endomorphism bundles
In this section we construct endomorphism bundles vector bundles constructed from line bundles. These will be the buildings blocks in the recursive subhomogeneous decomposition of the orbit-breaking algebra. We discuss the local trivialization of these bundles in detail.
First we recall that for a vector bundle $\mathscr{W}$ of rank $n$ over $X$, the endomorphism bundle $End(\mathscr{W})$ is a vector bundle over $X$ whose fibre at $x$ consists of all linear maps on the fibre vector space $\mathscr{W}_{x}$ of $\mathscr{W}$, namely $End(\mathscr{W})_{x}=End(\mathscr{W}_{x})$ for every $x$ (see [20, Example 6.8]).
Let $\mathcal{U}$ be an open cover of $X$ and let $\{(U,h_{U}:U\times\mathbb{C}^{n}\to\mathscr{W}|_{U})\}_{U\in\mathcal{U}}$ an atlas of $\mathscr{W}$ with transition functions $\{g_{U,V}\}_{U,V\in\mathcal{U}}$. Then it induces an atlas $\{(U,H_{U}:U\times M_{n}(\mathbb{C})\to End(\mathscr{W})|_{U})\}$ of $End(\mathscr{W})$ as follows:
$H_{U}(x,M):=h_{U}(x,\cdot)\circ M\circ h_{U}(x,\cdot)^{-1}\in End(\mathscr{W}_{x}),$ (9)
for $M\in M_{n}(\mathbb{C})$, which is represented in the following diagram with $h_{U,x}:=h_{U}(x,\cdot)$.
[…]
Moreover, whenever $U,V\in\mathcal{U}$ satisfy $U\cap V\neq\emptyset$, the map
$H_{V}^{-1}\circ H_{U}:(U\cap V)\times M_{n}(\mathbb{C})\to(U\cap V)\times M_{n}(\mathbb{C})$
is well defined and it satisfies
$H_{V}^{-1}\circ H_{U}(x,M)$ $=\left(x,(h_{V,x})^{-1}\circ h_{U,x}\circ M\circ(h_{U,x})^{-1}\circ h_{V,x}\right)$
$=\left(x,\mathrm{Ad}(g_{V,U}(x))M\right).$
6.0.1. Endomorphism bundles $\mathscr{M}^{(n)}$
Let $\mathscr{V}$ be a line bundle over $X$. Let $\mathscr{D}^{(m,n)}$ and $\mathscr{D}^{(n)}$ be vector bundles over $X$ given by
$\mathscr{D}^{(n)}$ $:=\mathscr{V}^{(0)}\oplus\mathscr{V}^{(1)}\oplus\mathscr{V}^{(2)}\oplus\cdots\oplus\mathscr{V}^{(n-1)},$
$\mathscr{D}^{(m,n)}$ $:=\mathscr{V}^{(m)}\oplus\mathscr{V}^{(m+1)}\oplus\mathscr{V}^{(m+2)}\oplus\cdots\oplus\mathscr{V}^{(m+n-1)},$
and let
$\mathscr{M}^{(n)}:=End(\mathscr{D}^{(n)}).$
Since $\mathscr{D}^{(n)}$ is of rank $n$, the bundle $\mathscr{M}^{(n)}$ has fiber $\mathscr{M}^{(n)}_{x}=End(\mathscr{D}^{(n)}_{x})$ isomorphic to $M_{n}(\mathbb{C})$.
6.0.2. Local trivialization of $\mathscr{M}^{(n)}$
We fix a finite open cover $\mathcal{U}_{0}$ of $X$ such that $\mathscr{V}$ is trivial over $\alpha^{i}(U)$ with a chart map
$h_{\alpha^{i}(U)}:\alpha^{i}(U)\times\mathbb{C}\to\mathscr{V}|_{\alpha^{i}(U)}$
for all $U\in\mathcal{U}_{0}$ and $0\leq i\leq 2n$. Then the atlas
$\{(\alpha^{i}(U),h_{\alpha^{i}(U)})\mid U\in\mathcal{U}_{0},\,i=0,\dots,2n\}$
gives us transition functions $g_{\alpha^{i}(U),\alpha^{j}(V)}:\alpha^{i}(U)\cap\alpha^{j}(V)\to U(1)$ such that
$h_{\alpha^{i}(U)}(x,\lambda)=h_{\alpha^{j}(V)}(x,g_{\alpha^{j}(V),\alpha^{i}(U)}(x)\lambda),\ \lambda\in\mathbb{C},$ (10)
for all $U\in\mathcal{U}_{0}$, $0\leq i\leq 2n$. Let
[…]
Then for the line bundle $\mathscr{V}^{(l)}$, $10\leq l\leq n-$ 1, the following maps
$(x,\lambda)\mapsto\lambda\,h_{\alpha^{l-1}(U)}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{\alpha(U)}(\alpha(x),1)\otimes h_{U}(x,1)$
from $U\times\mathbb{C}$ onto $\mathscr{V}^{(l)}|_{U}$, $U\in\mathcal{U}$, form an atlas. For convenience, we set
$v^{(l)}_{U}(x):=h_{\alpha^{l-1}(U)}(\alpha^{l-1}(x),1)\otimes\cdots\otimes h_{\alpha(U)}(\alpha(x),1)\otimes h_{U}(x,1),\ x\in U.$
Then $v_{U}^{(l)}:U\to\mathscr{V}^{(l)}|_{U}$ is a nonzero local section and thus each of the vectors $v^{(l)}_{U}(x)$, $U\in\mathcal{U}$ with $x\in U$, serves as a basis element of the one dimensional space $\mathscr{V}^{(l)}_{x}$. By (10) $v^{(l)}_{U}(x)$ and $v^{(l)}_{V}(x)$ satisfy the relation
$v^{(l)}_{U}(x)=\Pi_{i=0}^{l-1}g_{\alpha^{j}(V),\alpha^{j}(U)}(\alpha^{j}(x))\,v^{(l)}_{V}(x)$
whenever $x\in U\cap V$ for $U,V\in\mathcal{U}$. In turn, one has a basis of the $n$-dimensional space $\mathscr{D}^{(n)}_{x}=\mathscr{V}^{(0)}_{x}\oplus\mathscr{V}^{(1)}_{x}\oplus\cdots\oplus\mathscr{V}^{(n-1)}_{x}$ at each $x\in U$ given by
$e_{U,0}(x)$ $:=(v_{U}^{(0)}(x),0,0,\ldots,0),$
$e_{U,1}(x)$ $:=(0,v_{U}^{(1)}(x),0,\dots,0),$
[…]
$e_{U,n-1}(x)$ $:=(0,0,\dots,0,v_{U}^{(n-1)}(x))$
with convention $v_{U}^{(0)}(x):=1$. Then
$(\lambda_{0},\dots,\lambda_{n-1})\mapsto\lambda_{0}e_{U,0}(x)+\cdots+\lambda_{n-1}e_{U,n-1}(x):\mathbb{C}^{n}\to\mathscr{D}_{x}^{(n)},\ x\in U,$
gives us a chart map of $\mathscr{D}^{(n)}_{U}$ (by continuity of $e_{U,i}$’s over $U$) from which we obtain a chart map
$H_{U}^{(n)}:U\times M_{n}(\mathbb{C})\to\mathscr{M}_{U}^{(n)}:=\mathscr{M}^{(n)}|_{U}$
of the endomorphism bundle $\mathscr{M}^{(n)}=End(\mathscr{D}^{(n)})$ over $U\in\mathcal{U}$. Hence for a linear map $\varsigma(x)\in\mathscr{M}^{(n)}_{x}$, $x\in U$, there exists a unique matrix $\widetilde{\varsigma}_{U}(x)\in M_{n}(\mathbb{C})$ such that $H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x))=\varsigma(x)$. Namely, if $\widetilde{\varsigma}_{U}(x)$ is written as
$\widetilde{\varsigma}_{U}(x)=\footnotesize{\begin{bmatrix}a_{00}(x)&a_{01}(x)&a_{02}(x)&\cdots&a_{0(n-1)}(x)\\ a_{10}(x)&a_{11}(x)&a_{12}(x)&\cdots&a_{1(n-1)}(x)\\ a_{20}(x)&a_{21}(x)&a_{22}(x)&\cdots&a_{2(n-1)}(x)\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{(n-1)0}(x)&a_{(n-1)1}(x)&a_{(n-1)2}(x)&\cdots&a_{(n-1)(n-1)}(x)\end{bmatrix}},$
then it means that for each $i=0,\dots,n-1$,
$\varsigma(x)e_{U,i}(x)=a_{0i}(x)e_{U,0}(x)+a_{1i}(x)e_{U,1}(x)+\cdots+a_{(n-1)i}(x)e_{U,n-1}(x).$
When $x$ is fixed and clear in context, by abusing notation, we write simply $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$.
Now we consider what happens to the matrix representation of a linear map in $\mathscr{M}_{x}^{(n)}$ when we make a coordinate change in the vector space $\mathscr{D}^{(n)}_{x}$ for $x$ in an overlap $U\cap V$. For this, first set
$g_{VU}^{(l)}(x):=\prod_{i=0}^{l-1}{g_{\alpha^{i}(V),\alpha^{i}(U)}}(\alpha^{i}(x))$
and consider the following diagonal unitary matrix
$u_{VU}(x):=diag(1,g_{VU}^{(1)}(x),g_{VU}^{(2)}(x),\ldots,g_{VU}^{(n-1)}(x)).$ (11)
It is easily checked that
$\widetilde{\varsigma}_{V}(x)=u_{VU}(x)\widetilde{\varsigma}_{U}(x)u_{VU}^{-1}(x).$ (12)
In other words,
$H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x))=H_{V}^{(n)}(x,\mathrm{Ad}(u_{V,U}(x))\widetilde{\varsigma}_{U}(x)),$
and $x\mapsto\mathrm{Ad}(u_{V,U}(x)):X\to Aut(M_{n}(\mathbb{C}))$ are the transition functions for the atlas $\{(U,H_{U}^{(n)})\}_{U\in\mathcal{U}}$ of $\mathscr{M}^{(n)}$, where $\mathrm{Ad}(u)$ is an automorphism such that $\mathrm{Ad}(u)M=uMu^{*}$ for $M\in M_{n}(\mathbb{C})$.
Definition 6.1.
We call the atlas $\{(U,H_{U}^{(n)})\}_{U\in\mathcal{U}}$ the local trivialization induced from the atlas $\{h_{\alpha^{j}(U)}:\alpha^{j}(U)\times\mathbb{C}\to\mathscr{V}|_{\alpha^{j}(U)}\mid U\in\mathcal{U},\ j=0,\dots,n\}$.
If $\widetilde{\varsigma}_{U}(x)$ is written as above then with $(x)$ omitted,
$\widetilde{\varsigma}_{V}$ $=u_{VU}\widetilde{\varsigma}_{U}{u_{VU}}^{-1}$
$=\footnotesize{\begin{bmatrix}a_{00}&a_{01}{g_{VU}^{(1,0)}}^{-1}&a_{02}{g_{VU}^{(2,0)}}^{-1}&\cdots&a_{0(n-1)}{g_{VU}^{(n-1,0)}}^{-1}\\ g_{VU}^{(1,0)}a_{10}&a_{11}&a_{12}{g_{VU}^{(2,1)}}^{-1}&\cdots&a_{1(n-1)}{g_{VU}^{(n-1,1)}}^{-1}\\ g_{VU}^{(2,0)}a_{20}&{g_{VU}^{(2,1)}}a_{21}&a_{22}&\cdots&a_{2(n-1)}{g_{VU}^{(n-1,2)}}^{-1}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ g_{VU}^{(n-1,0)}a_{(n-1)0}&{g_{VU}^{(n-1,1)}}a_{(n-1)1}&{g_{VU}^{(n-1,2)}}a_{(n-1)2}&\cdots&a_{(n-1)(n-1)}\end{bmatrix}},$
where $g_{VU}^{(k,0)}(x):=g_{VU}^{(k)}(x)$ and $g_{VU}^{(k,j)}(x):=\prod_{i=j}^{k-1}{g_{\alpha^{i}(V)\alpha^{i}(U)}}(\alpha^{i}(x))$. Note that the matrices $\widetilde{\varsigma}_{U}(x)$ and $\widetilde{\varsigma}_{V}(x)$ have the same diagonal elements, that is, the diagonal is invariant under the coordinate change from $\{e_{U,i}(x)\}$ to $\{e_{V,i}(x)\}$ for $x\in U\cap V$. Also if one of the matrices is $m$th subdiagonal, then so is the other for $11\leq m\leq n-$ 1.
By a matrix representation of a linear map $\varsigma(x)$ in $\mathscr{M}^{(n)}_{x}$, we always mean a matrix $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ for some $U\in\mathcal{U}$ with $x\in U$ if not mentioned otherwise. We have the following remark from the above observations.
Remark 6.2.
Let $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and let $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ for $U\in\mathcal{U}$ and $x\in U$. Then we have the following:
  1. (1)
    The main diagonal is invariant: $\mathrm{diag}(\widetilde{\varsigma}_{U}(x))=\mathrm{diag}(\widetilde{\varsigma}_{V}(x))$ whenever $x\in U\cap V$.
  2. (2)
    If $\widetilde{\varsigma}_{U}(x)$ is $m$th subdiagonal, then so is $\widetilde{\varsigma}_{V}(x)$ whenever $x\in U\cap V$.
  3. (3)
    If $\varsigma(x)$ is represented as an $m$th (lower) subdiagonal matrix $\widetilde{\varsigma}_{U}(x)=(H_{U}^{(n)})^{-1}(\varsigma(x))$ on $U$ for $11\leq m\leq n-$ 1, then $\widetilde{\varsigma}_{U}(x)$ is of the form
[…]
    where each $s_{i,j}$ is a continuous function on $U$. In this case $\varsigma(x)$ maps each element $e_{U,i}(x)$ in basis of $\mathscr{D}^{(n)}_{x}$ to
    $\varsigma(x)e_{U,i}(x)=\begin{cases}s_{m+i,i}(x)e_{U,m+i}(x),&\text{for $0\leq i<n-m$},\\ 0,&\text{for $n-m\leq i\leq n-1$}.\end{cases}$ (13)
    We recall the definition of the basis element $e_{U,i}(x)$ and see that $\varsigma(x)$ maps each $\mathscr{V}^{(i)}_{x}$ to $\mathscr{V}^{(i+m)}_{x}$ for $0\leq i<n-m$, and furthermore, the linear map
    $\varsigma(x)|_{\mathscr{V}^{(i)}_{x}}:\mathscr{V}^{(i)}_{x}\to\mathscr{V}^{(m+i)}_{x}=\mathscr{V}_{\alpha^{i}(x)}^{(m)}\otimes\mathscr{V}^{(i)}_{x}$
    determines a vector $\varsigma^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that
    $\varsigma(x)|_{\mathscr{V}^{(i)}_{x}}:v\mapsto\varsigma^{(i)}(x)\otimes v,\quad v\in\mathscr{V}^{(i)}_{x},$ (14)
    since $\mathscr{V}^{(i)}_{x}$ and $\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ are all 1-dimensional. Using (13) we see that
    $\varsigma^{(i)}(x)=s_{m+i,i}(x)v^{(m)}_{\alpha^{i}(U)}(\alpha^{i}(x)),$ (15)
    since
    $v^{(m)}_{\alpha^{i}(U)}(\alpha^{i}(x))={h_{\alpha^{m-1}(\alpha^{i}(U))}}(\alpha^{m-1}(\alpha^{i}(x)),1)\otimes\cdots\otimes{h_{\alpha^{i}(U)}}(\alpha^{i}(x),1).$
Note that $\Gamma(\mathscr{M}^{(n)})$ is an algebra with respect to multiplication defined by pointwise composition of endomorphisms at each fibre. For $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and $x\in U\cap V$, we have
$H_{U}^{(n)}(x,(H_{U}^{(n)})^{-1}(\varsigma(x))^{*})$ $=\ H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x)^{*})$
$=\ H_{V}^{(n)}(x,u_{VU}\widetilde{\varsigma}_{U}(x)^{*}u_{VU}^{*})$
$=\ H_{V}^{(n)}(x,(u_{VU}\widetilde{\varsigma}_{U}(x)u_{VU})^{*})$
$=\ H_{V}^{(n)}(x,\widetilde{\varsigma}_{V}(x)^{*})$
$=\ H_{V}^{(n)}(x,(H_{V}^{(n)})^{-1}(\varsigma(x))^{*}),$
hence $\varsigma^{*}$ is well-defined by $\varsigma^{*}(x):=H_{U}^{(n)}(x,(H_{U}^{(n)})^{-1}(\varsigma(x))^{*})$. The -algebra $\Gamma(\mathscr{M}^{(n)})$ is in fact a $C^{*}$-algebra in the norm
$\|\varsigma\|:=\sup_{x\in X}\|(H_{U}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}.$
Note that $\|(H_{U}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}=\|(H_{V}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}$, so the norm above does not depend on the choice of the set $U$.
From now on, we will assume that we work with a fixed minimal homeomorphism $\alpha$ on a compact metric space $X$, a fixed line bundle $\mathscr{V}=[V,p,X]$ and a fixed closed subset $Y\subseteq X$ with non-empty interior.
The purpose of this paper is to show that the associated orbit-breaking algebra $\mathcal{O}(C_{0}(Y\setminus X)\Gamma(\mathscr{V}))$ has recursive subhomogeneous algebra structure. Since the proof is quite long, we will describe the main steps here.
We will denote $\Gamma(\mathscr{V},\alpha)$ by $\mathcal{E}$ and let $\mathcal{E}_{Y}$ be the orbit breaking bimodule $C_{0}(X\setminus Y)\mathcal{E}$.
Definition 6.3.
For $k=1,2,\ldots,K$, we define $\mathscr{M}_{k}$ to be the restriction of the endomorphism bundle $\mathscr{M}^{(r_{k})}$ onto $\overline{Y_{k}}$, namely
$\mathscr{M}_{k}:=\mathscr{M}^{(r_{k})}|_{\overline{Y_{k}}}=End(\mathscr{D}^{(r_{k})})|_{\overline{Y_{k}}}.$
For convenience, we write $H_{U}^{(k)}:=H_{U}^{(r_{k})}$, $U\in\mathcal{U}$, for the rest of the paper.
7. Embedding $\mathcal{O}(\mathcal{E}_{Y})$ into a direct sum of homogeneous algebras
In this section we show that there exists an embedding $\pi$ of the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$, where $\Gamma(\mathscr{M}_{k})$ is the $C^{*}$-algebra of continuous sections of the endomorphism bundle $\mathscr{M}_{k}$. We begin with constructing a $\mathrm{C}^{*}$-homomorphism $\pi_{k}$ from $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ for every $k=1,\dots,K$. Then we set $\pi=\pi_{1}\oplus\dots\oplus\pi_{K}$. We end this section with proving that $\pi_{1}$ is surjective while it is not true for $\pi_{k}$ when $2\leq k\leq K$ unless $Y_{k}$ is closed.
To construct a *-homomorphism $\pi_{k}$ from the Cuntz–Pimsner algebra $\mathcal{O}(\mathcal{E}_{Y})$ into the $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$, we define a covariant representation $(\pi_{k},\tau_{k})$ of $\mathcal{E}_{Y}$ into $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ which gives rise to the integrated -homomorphism of $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ which we will also denote by $\pi_{k}$.
Definition 7.1.
For $f\in C(X)$, $\xi\in\mathcal{E}_{Y}$, and $x\in\overline{Y_{k}}$, let $\pi_{k}(f)(x)$ and $\tau_{k}(\xi)(x)$ be the linear maps on $\mathscr{D}_{x}^{(r_{k})}$ given by
$\pi_{k}(f)(x)(\mathbf{a})$ $=(f(x)a_{0},f(\alpha(x))a_{1},\ldots,f(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$\tau_{k}(\xi)(x)(\mathbf{a}),$ $=(0,\xi(x)\otimes a_{0},\xi(\alpha(x))\otimes a_{1},\ldots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
[…]
Let $x\in Y_{k}\cap U$, $U\in\mathcal{U}$, and consider the basis $\{e_{U,0}(x),\dots,e_{U,r_{k}-1}(x)\}$ of $\mathscr{D}^{(r_{k})}_{x}$. If $\xi\in\mathcal{E}_{Y}$, then for each $j=0,\dots,r_{k}-2$, the scalar $\widetilde{\xi}(\alpha^{j}(x))\in\mathbb{C}$ with
$\xi(\alpha^{j}(x))=\widetilde{\xi}(\alpha^{j}(x))h_{\alpha^{j}(U)}(\alpha^{j}(x),1),$ (16)
satisfies
$\tau_{k}(\xi)(x)e_{U,j}(x)=\widetilde{\xi}(\alpha^{j}(x))e_{U,j+1}(x),$ (17)
and for the unitary $U_{x}:\mathbb{C}^{r_{k}}\to\mathscr{D}^{(r_{k})}_{x}$,
$U_{x}(\lambda_{0},\dots,\lambda_{r_{k}-1})=\lambda_{0}e_{U,0}(x)+\cdots+\lambda_{r_{k}-1}e_{U,r_{k}-1}(x),$
used above, we have the following matrix representation
$M_{\tau_{k}(\xi)}(x)=\begin{bmatrix}0&0&\cdots&0&0&0\\ \widetilde{\xi}(x)&0&\cdots&0&0&0\\ 0&\widetilde{\xi}(\alpha(x))&\ddots&0&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\cdots&\widetilde{\xi}(\alpha^{r_{k}-3}(x))&0&0\\ 0&0&\cdots&0&\widetilde{\xi}(\alpha^{r_{k}-2}(x))&0\end{bmatrix}$ (18)
of $U_{x}^{*}\tau_{k}(\xi)(x)U_{x}$, so
$\tau_{k}(\xi)(x)=U_{x}M_{\tau_{k}(\xi)}(x)U_{x}^{*}.$
This matrix representation will be used in the subsequent discussions along with the fact that the matrix depends on the choice of the open set $U$ containing $x$, but the matrix entries differ only by multiplication by scalars of modulus one. For $f\in C(X)$, the corresponding matrix representation is clearly
$M_{\pi_{k}(f)}(x)=\mathrm{diag}(f(x),f(\alpha(x)),\dots,f(\alpha^{r_{k}-1}(x))).$ (19)
It is rather obvious that $\pi_{k}$ is a -homomorphism, and we may write $\pi_{k}(f)(x)$ as
$\pi_{k}(f)(x)=\mathrm{diag}(f(x),f(\alpha(x)),\ldots,f(\alpha^{r_{k}-1}(x))).$
One can show that $\tau_{k}(\xi)^{*}\in\Gamma(\mathscr{M}_{k})$ is given by at $x\in\overline{Y_{k}}$,
$\tau_{k}(\xi)^{*}(x)(a_{0},\ldots,a_{r_{k}-1})=({\xi(x)}^{*}\otimes a_{1},\ldots,\xi(\alpha^{r_{k}-2}(x))^{*}\otimes a_{r_{k}-1},0),$ (20)
where $\xi(x)^{*}\otimes a_{1}:=\langle\xi(x),a_{1}\rangle_{\mathscr{V}_{x}}$ and
$\xi(\alpha^{i-1}(x))^{*}\otimes a_{i}:=\langle\xi(\alpha^{i-1}(x)),a^{i-1}_{i}\rangle_{\mathscr{V}_{\alpha^{i-1}(x)}}\,a_{i}^{i-2}\otimes\cdots\otimes a_{i}^{1}\otimes a_{i}^{0}\in\mathscr{V}_{x}^{(i-1)}$
for $a_{i}=a_{i}^{i-1}\otimes\cdots\otimes a_{i}^{1}\otimes a_{i}^{0}\in\mathscr{V}^{(i)}_{x}$, $12\leq i\leq r_{k}-$ 1.
Lemma 7.2.
Let $\mathcal{E}_{Y}:=C_{0}(X\setminus Y)\mathcal{E}$. Then we have the following.
  1. (a)
    $f\mapsto\pi_{k}(f):C(X)\to\Gamma(\mathscr{M}_{k})$ is a -homomorphism.
  2. (b)
    $\xi\mapsto\tau_{k}(\xi):\mathcal{E}_{Y}\to\Gamma(\mathscr{M}_{k})$ is a linear map.
  3. (c)
    $\pi_{k}(\langle\xi,\eta\rangle_{\mathcal{E}})=\tau_{k}(\xi)^{*}\tau_{k}(\eta)$, for $\xi,\eta\in\mathcal{E}_{Y}$.
  4. (d)
    $\tau_{k}(f\xi)=\pi_{k}(f)\tau_{k}(\xi)=\tau_{k}(\xi)\pi_{k}(f\circ\alpha)$.
  5. (e)
    $\pi_{k}(_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle)=\tau_{k}(\xi)\tau_{k}(\eta)^{*}$, where ${}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle:=\langle\eta,\xi\rangle_{\mathcal{E}}\circ\alpha^{-1}$ is the left inner product inherited from $\mathcal{E}_{Y}\subset\mathcal{E}$.
Proof.
(a) and (b) are clear.
(c) We have
$\pi_{k}(\langle\xi,\eta\rangle_{\mathcal{E}})(x)$ $=diag(\langle\xi,\eta\rangle_{\mathcal{E}}(x),\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha(x)),\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x))),$
and applying (20),
$\ \phantom{=}\tau_{k}(\xi)^{*}\tau_{k}(\eta)(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
$=\tau_{k}(\xi)^{*}(x)(0,\eta(x)\otimes a_{0},\eta(\alpha(x))\otimes a_{1},\ldots,\eta(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=(\xi(x)^{*}\otimes\eta(x)\otimes a_{0},\xi(\alpha(x))^{*}\otimes\eta(\alpha(x))\otimes a_{1},$
$\qquad\qquad\qquad\ldots,\xi(\alpha^{r_{k}-2}(x))\otimes\eta(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2},0)$
$=(\langle\xi,\eta\rangle_{\mathcal{E}}(x)a_{0},\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha(x))a_{1},\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-2}(x))a_{r_{k}-2},0)$
$=(\langle\xi,\eta\rangle_{\mathcal{E}}(x)a_{0},\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-2}(x))a_{r_{k}-2},\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$=\mathrm{diag}(\langle\xi,\eta\rangle_{\mathcal{E}}(x),\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha(x)),\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x)))(a_{0},a_{1},\ldots,a_{r_{k}-1}).$
since $\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x))a_{r_{k}-1}=0$.
(d) If $f\in C(X)$ and $\xi\in\mathcal{E}_{Y}$, then for any $(a_{0},\dots,a_{r_{k}-1})\in\mathscr{D}^{(r_{k})}_{x}$, we have
$\ (\pi_{k}(f)\tau_{k}(\xi))(x)(a_{0},\dots,a_{r_{k}-1})$
$=$ $\ \pi_{k}(f)(x)(0,\xi(x)\otimes a_{0},\dots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ (0,f(\alpha(x))\xi(x)\otimes a_{0},\dots,f(\alpha^{r_{k}-1}(x))\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ (0,(\xi f\circ\alpha)(x)\otimes a_{0},\dots,(\xi f\circ\alpha)(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ (0,(f\xi)(x)\otimes a_{0},\dots,(f\xi)(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ \tau_{k}(f\xi)(a_{0},\dots,a_{r_{k}-1}).$
(e) Since ${}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(x)=\langle\eta,\xi\rangle_{\mathcal{E}}\circ\alpha^{-1}(x)=0$ for $x\in\overline{Y_{k}}\subset Y$ and $\xi,\eta\in\mathcal{E}_{Y}$,
$\ \pi_{k}({}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle)(x)(a_{0},\dots,a_{r_{k}-1})$
$=$ $\ ({}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(x)a_{0},{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha(x))a_{1},\ldots,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$=$ $\ (0,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha(x)),\ldots,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha^{r_{k}-1}(x)))$
$=$ $\ (0,\langle\eta,\xi\rangle_{\mathcal{E}}(x)a_{1},\dots,\langle\eta,\xi\rangle_{\mathcal{E}}(\alpha^{i-1}(x))a_{i},\dots,\langle\eta,\xi\rangle_{\mathcal{E}}(\alpha^{r_{k}-2}(x))a_{r_{k}-1}).$
On the other hand, with $a_{i}=a_{i}^{\prime}\otimes a_{i}^{(i-1)}\in\mathscr{V}_{\alpha^{i-1}(x)}\otimes\mathscr{V}^{(i-1)}_{x}$, we have, for $11\leq i\leq r_{k}-$ 1,
$\ \tau_{k}(\xi)(x)\tau_{k}(\eta)^{*}(x)(a_{0},a_{1},\dots,a_{r_{k}-1})$
$=$ $\ \tau_{k}(\xi)(x)(\langle\eta(x),a_{1}\rangle_{\mathscr{V}_{x}},\ldots,\langle\eta(\alpha^{r_{k}-2}(x)),a_{r_{k}-1}^{\prime}\rangle_{\mathscr{V}_{\alpha^{r_{k}-2}(x)}}a_{r_{k}-1}^{(r_{k}-2)},0)$
$=$ $\ (0,\xi(x)\otimes\langle\eta(x),a_{1}\rangle_{\mathscr{V}_{x}},\dots$
$\dots,\xi(\alpha^{r_{k}-2}(x))\otimes\langle\eta(\alpha^{r_{k}-2}(x)),a_{r_{k}-1}^{\prime}\rangle_{\mathscr{V}_{\alpha^{r_{k}-2}(x)}}a_{r_{k}-1}^{(r_{k}-2)}).$
Thus it is enough to see that for each $11\leq i\leq r_{k}-$ 1,
$\langle\eta,\xi\rangle_{\mathcal{E}}(\alpha^{i-1}(x))a_{i}=\xi(\alpha^{i-1}(x))\otimes\langle\eta(\alpha^{i-1}(x)),a_{i}^{\prime}\rangle_{\mathscr{V}_{\alpha^{i-1}(x)}}a_{i}^{(i-1)}.$
But the left hand side
$\langle\eta,\xi\rangle_{\mathcal{E}}(\alpha^{i-1}(x))a_{i}=\langle\eta(\alpha^{i-1}(x)),\xi(\alpha^{i-1}(x))\rangle_{\mathscr{V}_{\alpha^{i-1}(x)}}a_{i}^{\prime}\otimes a_{i}^{(i-1)}$
must equal to the right hand side
$\xi(\alpha^{i-1}(x))\otimes\langle\eta(\alpha^{i-1}(x)),a_{i}^{\prime}\rangle_{\mathscr{V}_{\alpha^{i-1}(x)}}a_{i}^{(i-1)}$
$=\langle\eta(\alpha^{i-1}(x)),a_{i}^{\prime}\rangle_{\mathscr{V}_{\alpha^{i-1}(x)}}\xi(\alpha^{i-1}(x))\otimes a_{i}^{(i-1)}$
simply because of the fact that $\langle v_{1},v_{2}\rangle_{\mathscr{V}_{y}}v_{3}=\langle v_{1},v_{3}\rangle_{\mathscr{V}_{y}}v_{2}$ for any $y\in X$ and $v_{1},v_{2},v_{3}\in\mathscr{V}_{y}$ which can be seen by considering their inverse images under a chart map $h_{U}$ for $U\in\mathcal{U}$ with $y\in U$. ∎
Lemma 7.3.
  1. (a)
    The pair $(\pi_{k},\tau_{k})$ is a covariant representation of the $C(X)$-correspondence $\mathcal{E}_{Y}$, hence there exists a -homomorphism
    $\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$
    which we will denote by $\pi_{k}$ again.
  2. (b)
    There is a gauge action $\beta_{z}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k})$ for $z\in\mathbb{T}$.
Proof.
(a) Let $\varphi_{\mathcal{E}_{Y}}$ be the left action of $C(X)$ on $\mathcal{E}_{Y}$ which is the restriction of the left action of $C(X)$ to $\mathcal{E}$. Since $J_{\mathcal{E}_{Y}}:=\varphi_{\mathcal{E}_{Y}}^{-1}(\mathcal{K}(\mathcal{E}_{Y}))\cap(\ker\varphi_{\mathcal{E}_{Y}})^{\perp}=C_{0}(X\setminus Y)$ (this can be shown easily, we have to show that
$\pi_{k}(f)=\psi_{\tau_{k}}(\varphi_{\mathcal{E}_{Y}}(f)),\qquad f\in C_{0}(X\setminus Y),$
where $\psi_{\tau_{k}}:\mathcal{K}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ is the -homomorphism given by $\psi_{\tau_{k}}(\theta_{\xi,\eta})=\tau_{k}(\xi)\tau_{k}(\eta)^{*}$. But this follows from Lemma 7.2.(e) once we have the following
$\theta_{\xi,\eta}=\varphi_{\mathcal{E}_{Y}}(_{\mathcal{E}}\langle\xi,\eta\rangle),\qquad\xi,\eta\in\mathcal{E}_{Y}$
which actually holds since for any $\zeta\in\mathcal{E}_{Y}$,
$\theta_{\xi,\eta}(\zeta)=\xi\langle\eta,\zeta\rangle_{\mathcal{E}}={}_{\mathcal{E}}\langle\xi,\eta\rangle\zeta=\varphi_{\mathcal{E}_{Y}}(_{\mathcal{E}}\langle\xi,\eta\rangle)\zeta.$
Now by the universal property of $\mathcal{O}(\mathcal{E}_{Y})$, there exists a -homomorphism of $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ which we also denote $\pi_{k}$.
(b) For $z\in\mathbb{T}$, let $\beta_{z}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k})$ be the *-homomorphism given by $\beta_{z}=Ad(U_{z})$ on each fiber at $x$ for a unitary map $U_{z}$ which is an endomorphism on each fiber at $x$ by
$U_{z}:(a_{0},a_{1},a_{2},\ldots,a_{r_{k}-1})\mapsto(a_{0},za_{1},z^{2}a_{2},\ldots,z^{r_{k}-1}a_{r_{k}-1}).$
We see that for $f\in C(X)$,
$\ U_{z}\pi_{k}(f)(x)U_{z}^{*}(a_{0},a_{1},\,\ldots,\,a_{r_{k}-1})$
$=$ $\ U_{z}\pi_{k}(f)(x)(a_{0},\overline{z}a_{1},\overline{z}^{2}a_{2},\,\ldots,\,\overline{z}^{r_{k}-1}a_{r_{k}-1})$
$=$ $\ U_{z}(f(x)a_{0},f(\alpha(x))\overline{z}a_{1},f(\alpha^{2}(x))\overline{z}^{2}a_{2},\,\ldots,\,f(\alpha^{r_{k}-1}(x))\overline{z}^{r_{k}-1}a_{r_{k}-1})$
$=$ $\ (f(x)a_{0},zf(\alpha(x))\overline{z}a_{1},z^{2}f(\alpha^{2}(x))\overline{z}^{2}a_{2},\,\ldots,\,z^{r_{k}-1}f(\alpha^{r_{k}-1}(x))\overline{z}^{r_{k}-1}a_{r_{k}-1})$
$=$ $\ (f(x)a_{0},f(\alpha(x))a_{1},f(\alpha^{2}(x))a_{2},\,\ldots,\,f(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$=$ $\ \pi_{k}(f)(x)$
and for $\xi\in\mathcal{E}_{Y}$,
$\ U_{z}\tau_{k}(\xi)(x)U_{z}^{*}(a_{0},a_{1},\,\ldots,\,a_{r_{k}-1})$
$=$ $\ U_{z}\tau_{k}(\xi)(x)(a_{0},\overline{z}a_{1},\overline{z}^{2}a_{2},\,\ldots,\,\overline{z}^{r_{k}-1}a_{r_{k}-1})$
$=$ $\ U_{z}(0,\xi(x)\otimes a_{0},\xi(\alpha(x))\otimes\overline{z}a_{1},\,\ldots,\,\xi(\alpha^{r_{k}-2}(x))\otimes\overline{z}^{r_{k}-2}a_{r_{k}-2})$
$=$ $\ (0,z\xi(x)\otimes a_{0},z^{2}\xi(\alpha(x))\otimes\overline{z}a_{1},\,\ldots,\,z^{r_{k}-1}\xi(\alpha^{r_{k}-2}(x))\otimes\overline{z}^{r_{k}-2}a_{r_{k}-2})$
$=$ $\ (0,z\xi(x)\otimes a_{0},\,\ldots,\,z\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ z\tau_{k}(\xi)(x).$
Thus the representation $(\pi_{k},\tau_{k})$ admits a gauge action for each $z\in\mathbb{T}$. ∎
Proposition 7.4.
Let $\pi_{k}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ be the -homomorphism obtained in Lemma 7.3 for $1\leq k\leq K$. Then
$\pi:=\oplus_{k=1}^{K}\pi_{k}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$
is an injective -homomorphism.
Proof.
To see that $\pi$ is injective, we only need to show that it is injective on $C(X)$ by [23, Theorem 6.4] since $\pi(\mathcal{O}(\mathcal{E}_{Y}))$ admits the gauge action
$z\mapsto\beta_{z}:=\beta_{1,z}\oplus\cdots\oplus\beta_{K,z},$
where $z\mapsto\beta_{k,z}$ is the gauge action on $\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$ obtained in Lemma 7.3. Let $f\in C(X)$ be a function such that $\pi(f)=\pi_{1}(f)\oplus\cdots\oplus\pi_{K}(f)=0$. Then
$\pi_{k}(f)(x)=\mathrm{diag}(f(x),f(\alpha(x)),\dots,f(\alpha^{r_{k}-1}(x))=(0,\dots,0)$
for every $x\in\overline{Y_{k}}$ and $1\leq k\leq K$. Sine $X=\cup_{k=1}^{K}\cup_{j=1}^{r_{k}-1}\alpha^{j}(\overline{Y_{k}})$ by (7), we have $f=0$ as desired. ∎
Remark 7.5.
If $a=\xi_{1}\xi_{2}\cdots\xi_{k}\xi_{k+1}^{*}\cdots\xi_{k+l}^{*}\in\mathcal{O}(\mathcal{E}_{Y})$ for $\xi_{i}\in\mathcal{E}_{Y}$, and $k+l$ is the minimal length for any possible such expressions of $a$, then we say that $\deg(a)=k-l$. Then the span of the elements of degree $n$ for $n\in\mathbb{Z}$ is dense in $\mathcal{O}(\mathcal{E}_{Y})$, namely
$\mathcal{O}(\mathcal{E}_{Y})=\overline{\oplus_{n\in\mathbb{Z}}\overline{E_{n}}},$
where $E_{n}$ denotes the $C(X)$-linear span of the elements of degree $n$. Note that $E_{0}=C(X)$, and for $n\geq 1$
$\xi_{1}\otimes\cdots\otimes\xi_{n}\mapsto\xi_{1}\cdots\xi_{n}$
extends to an isometric linear map of $\mathcal{E}_{Y}^{\otimes n}$ onto $E_{n}$ (see Remark 4.1 of [2]). Thus we may identify $\mathcal{E}_{Y}^{\otimes n}$ with the subspace $E_{n}$ of $\mathcal{O}(\mathcal{E}_{Y})$.
Let $\xi=\xi_{1}\xi_{2}\in E_{2}$ for $\xi_{1},\xi_{2}\in\mathcal{E}_{Y}$. Then for $(a_{0},a_{1},\ldots,a_{r_{k}-1})\in\mathscr{V}^{(0)}_{x}\oplus\mathscr{V}^{(1)}_{x}\oplus\mathscr{V}^{(2)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{k}-1)}_{x}$ and $x\in\overline{Y_{k}}$, since $\pi_{k}$ is a homomorphism, we have
$\pi_{k}(\xi_{1}\xi_{2})(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
=\ = $\pi_{k}(\xi_{1})(x)\pi_{k}(\xi_{2})(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
=\ = $\pi_{k}(\xi_{1})(x)(0,\xi_{2}(x)\otimes a_{0},\dots,\xi_{2}(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
=\ = $(0,0,\xi_{1}(\alpha(x))\otimes\xi_{2}(x)\otimes a_{0},\dots,\xi_{1}(\alpha^{r_{k}-2}(x))\otimes\xi_{2}(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3})$
=\ = $(0,0,\psi(\xi_{1}\otimes\xi_{2})(x)\otimes a_{0},\dots,\psi(\xi_{1}\otimes\xi_{2})(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3}).$
From a similar calculation and Proposition 3.1, we see that for $\xi\in\mathcal{E}_{Y}^{\otimes n}=E_{n}$, $\pi_{k}(\xi)(x)$ acts on $\,\mathscr{D}_{x}^{(r_{k})}$ as follows:
$\ \pi_{k}(\xi)(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
$=$ $\ (0,\ldots,0,\psi(\xi)(x)\otimes a_{0},\psi(\xi)(\alpha(x))\otimes a_{1},\ldots,\psi(\xi)(\alpha^{r_{k}-1-n}(x))\otimes a_{r_{k}-1-n})$
with zeros in the first $n$ coordinates, or on each component $\mathscr{V}_{x}^{(i)}$,
$\pi_{k}(\xi)(x)|_{\mathscr{V}_{x}^{(i)}}(a_{i})=\psi(\xi)(\alpha^{i}(x))\otimes a_{i}\in\mathscr{V}_{\alpha^{i}(x)}^{(n)}\otimes\mathscr{V}_{x}^{(i)}=\mathscr{V}_{x}^{(n+i)}.$ (21)
Thus for $k=1,\dots,K$,
$\pi_{k}(E_{n})=0\ \text{ whenever }n\geq r_{k}.$
Definition 7.6.
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. For an integer $10\leq m\leq r_{K}-$ 1, we say that $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th subdiagonal matrix representation if at every $x\in\overline{Y_{k}}$, the matrix $(H_{U}^{(k)})^{-1}(x,\varsigma_{k}(x))$ is $m$th subdiagonal for some (hence for all) $U\in\mathcal{U}$ with $x\in U$. Here $(H_{U}^{(k)})^{-1}(x,\varsigma_{k}(x))$ is regarded as the zero matrix if $m\geq r_{k}$.
Assume that each $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th lower subdiagonal matrix representation. We know from Remark 6.2 that for $x\in\overline{Y_{k}}$, the map
$\varsigma_{k}(x):\mathscr{D}^{(r_{k})}_{x}\to\mathscr{D}^{(r_{k})}_{x}$
is $m$th lower subdiagonal in matrix representation if and only if $\varsigma_{k}(x)$ maps $\mathscr{V}_{x}^{(i)}$ in $\mathscr{D}^{(r_{k})}_{x}=\mathscr{V}_{x}^{(0)}\oplus\mathscr{V}_{x}^{(1)}\oplus\cdots\oplus\mathscr{V}_{x}^{(r_{k}-1)}$ to $\mathscr{V}_{x}^{(m+i)}$ for $i<r_{k}-m$ and $\{0\}$ for $r_{k}-m\leq i<r_{k}$. Furthermore, the linear map
$\varsigma_{k}(x)|_{\mathscr{V}^{(i)}_{x}}:\mathscr{V}^{(i)}_{x}\to\mathscr{V}^{(m+i)}_{x}=\mathscr{V}_{\alpha^{i}(x)}^{(m)}\otimes\mathscr{V}^{(i)}_{x}$
is determined by a vector $\varsigma_{k}^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that
$\varsigma_{k}(x)|_{\mathscr{V}^{(i)}_{x}}:v\mapsto\varsigma_{k}^{(i)}(x)\otimes v,\ \text{for any $v\in\mathscr{V}^{(i)}_{x}$}.$ (22)
From the definition of $\pi_{k}$, we know that if $g\in C(X)$, then $\pi_{k}(g)(x)$ is represented as a diagonal matrix for all $x\in\overline{Y_{k}}$. Also if $\xi\in\mathcal{E}_{Y}^{\otimes m}\subset\mathcal{O}(\mathcal{E}_{Y})$ for some $11\leq m\leq r_{k}-$ 1, then $\pi_{k}(\xi)(x)$ is represented as an $m$th lower subdiagonal matrix since it maps $\mathscr{V}_{x}^{(i)}$ into $\mathscr{V}_{x}^{(i+m)}$ by (21). Also if $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})=\pi(\xi)$ for $\xi\in\mathcal{E}^{\otimes m}$, then we have $\varsigma_{k}^{(i)}(x)=\psi(\xi)(\alpha^{i}(x))$ by (21) and (14).
Lemma 7.7.
If $Y_{k}$ is closed, then $\pi_{k}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ is surjective. In particular, $\pi_{1}$ is surjective.
Proof.
If $k=1$ and $r_{1}=1$, then $\mathscr{D}^{(r_{1})}=\mathscr{V}^{(0)}=X\times\mathbb{C}$ is a trivial line bundle, hence $\mathscr{M}_{1}=\mathscr{M}^{(r_{1})}|_{Y_{1}}=Y_{1}\times M_{1}(\mathbb{C})=Y_{1}\times\mathbb{C}$. Thus $\Gamma(\mathscr{M}_{1})=C(Y_{1})$ and $\Gamma(\mathscr{M}_{1})=\pi_{1}(C(X))$ by definition of $\pi_{1}$.
We assume $r_{k}>1$. Since the $C^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ is generated by the sections $\varsigma\in\Gamma(\mathscr{M}_{k})$ which is an $m$th lower subdiagonal matrix in its matrix representation for $0\leq m<r_{k}$, it is enough to show that $\varsigma\in\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$ for any such section $\varsigma$. Note that if $0\leq i\neq j<r_{k}$, then $\alpha^{i}(Y_{k})\cap\alpha^{j}(Y_{k})=\emptyset$ by definition of $r_{k}$.
Let $\varsigma\in\Gamma(\mathscr{M}_{k})$ have $m$th lower subdiagonal matrix representation. Then we know from the discussion given before this lemma that for $i=0,\dots,r_{k}-m-1$, there exists $\varsigma^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that the map
$\varsigma(x)|_{\mathscr{V}^{(i)}_{x}}:\mathscr{V}^{(i)}_{x}\to\mathscr{V}^{(m+i)}_{x}=\mathscr{V}_{\alpha^{i}(x)}^{(m)}\otimes\mathscr{V}^{(i)}_{x}$
is given by $v\mapsto\varsigma^{(i)}(x)\otimes v$ for $v\in\mathscr{V}^{(i)}_{x}$.
Define $\xi\in\Gamma(\mathscr{V}^{(m)}|_{\cup_{i=0}^{r_{k}-m-1}\alpha^{i}(Y_{k})})$ on the closed subset $\cup_{i=0}^{r_{k}-m-1}\alpha^{i}(Y_{k})$ of $X$ by
$\xi|_{\alpha^{i}(Y_{k})}(\alpha^{i}(x)):=\varsigma^{(i)}(x)$
for $x\in Y_{k}$ and $i=0,\dots,r_{k}-m-1$. This $\xi$ is well-defined since the sets $\alpha^{i}(Y_{k})$ are mutually disjoint. It is easy to see that
$(\cup_{l=1}^{m}\alpha^{-l}(Y))\cap(\cup_{i=0}^{r_{k}-m-1}\alpha^{i}(Y_{k}))=\emptyset,$
and then by the Tietze extension theorem for a vector bundle (see [5, Section 1.4]), $\xi$ extends to a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over the whole space $X$ such that
$\widetilde{\xi}|_{\alpha^{i}(Y_{k})}=\xi|_{\alpha^{i}(Y_{k})}\ \ \text{and }\ \widetilde{\xi}|_{\cup_{l=1}^{m}\alpha^{-l}(Y)}=0.$
Recall from Proposition 3.2 that there is an isomorphism $\psi:\mathcal{E}^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$ such that the property $\widetilde{\xi}|_{\cup_{l=1}^{m}\alpha^{-l}(Y)}=0$ implies $\psi^{-1}(\widetilde{\xi})\in\mathcal{E}_{Y}^{\otimes m}=E_{m}\subset\mathcal{O}(\mathcal{E}_{Y})$ (see Remark 7.5). Now for any $(a_{0},a_{1},\ldots,a_{r_{k}-1})\in\mathscr{D}^{(r_{k})}_{x}=\mathscr{V}^{(0)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{k}-1)}_{x}$, we then have
$\ \pi_{k}(\psi^{-1}(\widetilde{\xi}))(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
$=$ $\ (0,\ldots,0,\psi(\psi^{-1}(\widetilde{\xi}))(x)\otimes a_{0},\ldots,\psi(\psi^{-1}(\widetilde{\xi}))(\alpha^{r_{k}-m-1}(x))\otimes a_{r_{k}-m-1})$
$=$ $\ (0,\ldots,0,\widetilde{\xi}(x)\otimes a_{0},\widetilde{\xi}(\alpha(x))\otimes a_{1},\ldots,\widetilde{\xi}(\alpha^{r_{k}-m-1}(x))\otimes a_{r_{k}-m-1})$
$=$ $\ (0,\ldots,0,\varsigma^{(0)}(x)\otimes a_{0},\varsigma^{(1)}(x)\otimes a_{1},\ldots,\varsigma^{(r_{k}-m-1)}(x)\otimes a_{r_{k}-m-1})$
$=$ $\ \varsigma(x)(a_{0},a_{1},\ldots,a_{r_{k}-1}),$
hence $\pi_{k}(\psi^{-1}(\widetilde{\xi}))=\varsigma$ follows. ∎
8. A recursive subhomogeneous decomposition of $\mathcal{O}(\mathcal{E}_{Y})$
In this section we explicitly describe the image of $\mathcal{O}(\mathcal{E}_{Y})$ under the injective -homomorphism $\pi:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. In particular, we construct a recursive subhomogeneous decomposition for the range of $\pi$ which yields that the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ is a recursive subhomogeneous algebra. As in the case of $\mathrm{C}^{*}$-algebras of minimal homeomorphisms, we investigate the behaviour of boundary points. However, the presence of a twist (that is, a non-trivial line bundle) leads to further difficulties since we need to consider the local trivialization of the underlying endomorphism bundles when we are tracking the travel of boundary points through the Rokhlin towers.
Note that the sets $Y_{k}$ are in general not closed. If $x\in\overline{Y_{k}}$, by continuity of $\alpha$ and the fact that $Y$ is closed, we have $\alpha^{r_{k}}(x)\in Y$. Whereas, if $x\in\overline{Y_{k}}\setminus Y_{k}$, then this will not be the first return time of $x$ so that $x\in Y_{t_{1}}$ for some $t_{1}<k$, and hence $\alpha^{r_{t_{1}}}(x)\in Y_{t_{2}}$ for some $t_{2}<k$. It follows that $x\in\overline{Y_{k}}\cap Y_{t_{1}}\cap\alpha^{-{r_{t_{1}}}}(Y_{t_{2}})$. If $r_{t_{1}}+r_{t_{2}}<r_{k}$, then there is some $t_{3}<k$ such that $\alpha^{r_{t_{1}}+r_{t_{2}}}(x)\in Y_{t_{3}}$. Continuing in this manner, we get
$t_{1},t_{2},\ldots,t_{m(x,k)}<k$
such that
$r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)}}=r_{k}$
and
$x\in\overline{Y_{k}}\cap Y_{t_{1}}\cap\alpha^{-{r_{t_{1}}}}(Y_{t_{2}})\cap\cdots\cap\alpha^{-(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)-1}})}(Y_{t_{m(x,k)}})$
Define $\mu_{x,k}$ to be the ordered multi-set
$\mu_{x,k}:=\{t_{1},t_{2},\ldots,t_{m(x,k)}\},$
and for $1\leq s\leq m(x,k)$ let $\mu_{x,k}(s)=t_{s}$ and $|\mu_{x,k}|=m(x,k)$. Let
[…]
for $x\in\overline{Y_{k}}$, $2\leq k\leq K$, Note that the vector space
$\mathscr{D}^{(r_{k})}_{x}$ $=\mathscr{V}^{(0)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{t_{1}}-1)}_{x}\oplus\mathscr{V}^{(r_{t_{1}})}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{t_{1}}+r_{t_{2}}-1)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{k}-1)}_{x}$
$=\mathscr{D}_{x}^{(r_{t_{1}})}\oplus\mathscr{D}_{x}^{(R_{x,k,1},r_{t_{2}})}\oplus\mathscr{D}_{x}^{(R_{x,k,2},r_{t_{3}})}\oplus\cdots\oplus\mathscr{D}_{x}^{(R_{x,k,m(x,k)-1},r_{t_{m(x,k)}})}$
is the direct sum of $\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}$, $s=1,2,\ldots,m(x,k)$, where $\mathscr{D}^{(0,r_{t_{1}})}:=\mathscr{D}^{(r_{t_{1}})}$. We call
$\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}=\mathscr{V}_{x}^{(R_{x,k,s-1})}\oplus\mathscr{V}_{x}^{(R_{x,k,s-1}+1)}\oplus\cdots\oplus\mathscr{V}_{x}^{(R_{x,k,s-1}+r_{t_{s}}-1)}$
the $s$th component of $\mathscr{D}_{x}^{(r_{k})}$ for $s=1,\dots,m(x,k)$, which can also be written as follows:
$\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}$
$=\ \mathscr{V}_{x}^{(R_{x,k,s-1})}\oplus\mathscr{V}_{x}^{(R_{x,k,s-1}+1)}\oplus\cdots\oplus\mathscr{V}_{x}^{(R_{x,k,s}-1)}$
$=\ \mathscr{V}_{x}^{(R_{x,k,s-1})}\oplus\mathscr{V}_{x}^{(R_{x,k,s-1}+1)}\oplus\cdots\oplus\mathscr{V}_{x}^{(R_{x,k,s-1}+r_{t_{s}}-1)}$
$=\ \big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}\big{)}\oplus\cdots\oplus\big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}\big{)}$
$=\ \big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\oplus\cdots\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}\big{)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}$
$=\ \mathscr{D}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}})}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}.$
The next definition is based on [27, Lemma 4].
Definition 8.1.
For each $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$, we say that $\varsigma$ has the boundary decomposition property if it satisfies the following: for any $x\in\overline{Y_{k}}\setminus Y_{k}$, $k=1,\dots,K$ with
$x\in\overline{Y_{k}}\cap Y_{t_{1}}\cap\alpha^{-{r_{t_{1}}}}(Y_{t_{2}})\cap\cdots\cap\alpha^{-(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)-1}})}(Y_{t_{m(x,k)}})$
and $r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)}}=r_{k}$, the linear map $\varsigma_{k}(x)\in\mathscr{M}_{k}|_{x}=End(\mathscr{D}_{x}^{(r_{k})})$ acts on each $s$th component
$\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}=\ \mathscr{D}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}})}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}$
of $\mathscr{D}^{(r_{k})}_{x}$ as
$\varsigma_{k}(x)|_{\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}}=\varsigma_{t_{s}}(\alpha^{R_{x,k,s-1}}(x))\otimes id_{\mathscr{V}^{(R_{x,k,s-1})}_{x}},$ (23)
or equivalently, $\varsigma_{k}(x)$ has its matrix representation of the following form
$\widetilde{\varsigma}_{k}(x)=\begin{bmatrix}\widetilde{\varsigma}_{t_{1}}(x)&\\ &\widetilde{\varsigma}_{t_{2}}(\alpha^{R_{x,k,1}}(x))&\\ &&\ddots\\ &&&\widetilde{\varsigma}_{t_{m(x,k)}}(\alpha^{R_{x,k,m(x,k)-1}}(x))\end{bmatrix}$
[…]
Next we show that the image of $\pi$ can be determined by the boundary decomposition property.
Lemma 8.2.
If $\varsigma\in\pi(\mathcal{O}(\mathcal{E}_{Y}))$, then $\varsigma$ has the boundary decomposition property.
Proof.
Since it is not hard to see that the boundary decomposition property is preserved under addition, multiplication, and involution, we only need to show the assertion when $\varsigma=\pi(g)$ for $g\in C(X)$ or $\varsigma=\pi(\xi)$ for $\xi\in\mathcal{E}_{Y}$.
First let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})=\pi(g)\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ for some $g\in C(X)$ and let $x\in\overline{Y_{k}}\setminus Y_{k}$ be a boundary point of $Y_{k}$ as in Definition 8.1. We then have
$\varsigma_{k}(x)$ $=\pi_{k}(g)(x)=\mathrm{diag}(g(x),g(\alpha(x)),\ldots,g(\alpha^{r_{k}-1}(x))):\mathscr{D}^{(r_{k})}_{x}\to\mathscr{D}^{(r_{k})}_{x}$
and the restriction $\varsigma_{k}(x)|_{\mathscr{V}_{x}^{(i)}}:\mathscr{V}_{x}^{(i)}\to\mathscr{V}_{x}^{(i)}$ is a multiplication by a number which we write as $\varsigma_{k}^{(i)}(x)$ as observed in (14) with $m=0$. On $s$th component
$\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}=\big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(1)}\oplus\cdots\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}\big{)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})},$
of $\mathscr{D}^{(r_{k})}_{x}$, the map $\varsigma_{k}(x)=\pi_{k}(g)(x)$ acts as
$\mathrm{diag}(g(\alpha^{R_{x,k,s-1}}(x)),g(\alpha^{R_{x,k,s-1}+1}(x)),\ldots,g(\alpha^{R_{x,k,s-1}+r_{t_{s}}-1}(x))).$
Whereas $\varsigma_{t_{s}}\in\Gamma(\mathscr{M}_{t_{s}})$ acts at $\alpha^{R_{x,k,s-1}}(x)\in Y_{t_{s}}$ as
$\ \varsigma_{t_{s}}(\alpha^{R_{x,k,s-1}}(x))$
$=$ $\ \pi_{t_{s}}(g)(\alpha^{R_{x,k,s-1}}(x))$
$=$ $\ \mathrm{diag}(g(\alpha^{R_{x,k,s-1}}(x)),g(\alpha(\alpha^{R_{x,k,s-1}}(x))),\ldots,g(\alpha^{r_{t_{s}}-1}(\alpha^{R_{x,k,s-1}}(x))))$
on the vector space
$\mathscr{D}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}})}$
$=\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(1)}\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(2)}\oplus\cdots\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}.$
Therefore on $s$th component $\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}=\mathscr{D}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}})}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}$,
$\varsigma_{k}(x)=\varsigma_{t_{s}}(\alpha^{R_{x,k,s-1}}(x))\otimes\mathrm{id}_{\mathscr{V}^{(R_{x,k,s-1})}_{x}},$
from which we have
$\varsigma_{k}^{(R_{x,k,s-1}+i)}(x)=\varsigma_{t_{s}}^{(i)}(\alpha^{R_{x,k,s-1}}(x)).$
Now let $\varsigma=\varsigma_{1}\oplus\cdots\oplus\varsigma_{K}=\pi_{1}(\xi)\oplus\cdots\oplus\pi_{K}(\xi)$ for some $\xi\in\mathcal{E}_{Y}$. Then $\pi_{k}(\xi)(x)$ acts on the $s$th component of $\mathscr{D}_{x}^{(r_{k})}$ as follows. If we have $(a_{0},a_{1},\dots,a_{r_{k}-1})\in\mathscr{D}_{x}^{(r_{k})}$, then its $s$th component is equal to
$(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})\in\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}$
and
$\varsigma_{k}(x)|_{\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}}(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
=\ = $\pi_{k}(\xi)(x)|_{\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}}(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
=\ = $(0,\xi(\alpha^{R_{x,k,s-1}}(x))\otimes a_{R_{x,k,s-1}},\dots)$
=\ = $\pi_{t_{s}}(\xi)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
=\ = $\varsigma_{t_{s}}(x)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1}),$
which proves the lemma. ∎
Recall from (14) that we have a map
$x\mapsto\varsigma_{k}^{(i)}(x):\overline{Y_{k}}\to\mathscr{V}_{\alpha^{i}(x)}^{(m)}$
obtained by restricting the linear map $\varsigma_{k}(x):\mathscr{D}^{(r_{k})}_{x}\to\mathscr{D}^{(r_{k})}_{x}$ onto $\mathscr{V}_{x}^{(i)}$, a direct summand of $\mathscr{D}^{(r_{k})}_{x}$.
Lemma 8.3.
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ have the boundary decomposition property, and let $0\leq m<r_{K}$. If $\varsigma_{k}$ has an $m$th subdiagonal matrix representation for all $k$, then there is a section $G\in\Gamma(\mathscr{V}^{(m)})$ of $\mathscr{V}^{(m)}$ such that
$G\big{|}_{\alpha^{i}(\overline{Y_{k}})}=\varsigma_{k}^{(i)}\circ\alpha^{-i}\big{|}_{\alpha^{i}(\overline{Y_{k}})},$
for all $0\leq i\leq r_{k}-1,$ and vanishes on $\alpha^{-m}(Y)\cup\alpha^{-m+1}(Y)\cup\cdots\cup\alpha^{-1}(Y)$.
Proof.
To construct a continuous section $G:X\to\mathscr{V}^{(m)}$ satisfying the properties stated in this lemma, we first note $X=\cup_{k=1}^{K}\cup_{i=0}^{r_{k}-1}\alpha^{i}(\overline{Y_{k}})$ and define $G$ on each $\alpha^{i}(\overline{Y_{k}})$ in such a way that the values of $G$ match up at any possible overlaps $\alpha^{j}(\overline{Y_{l}})\cap\alpha^{i}(\overline{Y_{k}})$.
Let $x\in\overline{Y_{k}}\setminus Y_{k}$ be a point as in Definition 8.1. Since $\varsigma$ has the boundary decomposition property, on each $s$th component $\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}$ of $\mathscr{D}^{(r_{k})}_{x}$,
$\varsigma_{k}(x)|_{\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}}=\varsigma_{t_{s}}(\alpha^{R_{x,k,s-1}}(x))\otimes id_{\mathscr{V}^{(R_{x,k,s-1})}_{x}},$ (24)
for $s=2,\dots,m(x,k)$. Also by assumption, $\varsigma_{k}(x):\mathscr{D}^{(r_{k})}_{x}\to\mathscr{D}^{(r_{k})}_{x}$ has an $m$th subdiagonal matrix representation, which means that $\varsigma_{k}(x)$ maps $\mathscr{V}_{x}^{(i)}$ into $\mathscr{V}_{x}^{(i+m)}=\mathscr{V}_{\alpha^{i}(x)}^{(m)}\otimes\mathscr{V}_{x}^{(i)}$. Hence, there exist elements $\varsigma_{k}^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that
$\varsigma_{k}(x)|_{\mathscr{V}_{x}^{(i)}}:v\mapsto\varsigma_{k}^{(i)}(x)\otimes v,\quad v\in\mathscr{V}^{(i)}_{x},$
where $\varsigma_{k}^{(i)}(x)=0$ for $i\geq r_{k}-m$. The equation (24) implies that for $i$ with $R_{x,k,s-1}\leq i<R_{x,k,s}$,
$\varsigma_{k}^{(i)}(x)=\varsigma_{t_{s}}^{(i-R_{x,k,s-1})}(\alpha^{R_{x,k,s-1}}(x)).$
Now we define $G:X\to\mathscr{V}^{(m)}$ by
$G\big{|}_{\alpha^{i}(\overline{Y_{k}})}=\varsigma_{k}^{(i)}\circ\alpha^{-i}\big{|}_{\alpha^{i}(\overline{Y_{k}})},$
for $i=0,1,2,\ldots,r_{k}-1$. $G$ is well-defined if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})$ implies
$\varsigma_{k}^{(i)}\circ\alpha^{-i}(x)=\varsigma_{l}^{(j)}\circ\alpha^{-j}(x).$ (25)
By (7), it is enough to show that if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})\cap\alpha^{p}(Y_{q})$, then (25) holds. Note that
$\alpha^{i}(\overline{Y_{k}})\cap\alpha^{p}(Y_{q})\neq\emptyset\ \text{ implies}\ i\geq p$ (26)
whenever $0\leq i<r_{k}$ and $0\leq p<r_{q}$. Since otherwise, $0<p-i<r_{q}$, but we would have $\overline{Y_{k}}\cap\alpha^{p-i}(Y_{q})\neq\emptyset$, contradicting with the fact that $Y\cap\alpha^{p-i}(Y_{q})=\emptyset$.
If $x=\alpha^{i}(y)\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{p}(Y_{q})$, for $p\leq i<r_{k}$ and $q$, which implies that $y\in\overline{Y_{k}}\setminus Y_{k}$, then
$\alpha^{i-p}(y)=\alpha^{-p}(x)\in Y_{q}.$
Since $0\leq i-p<r_{k}$, there is an $n\geq 0$ such that
$r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}}\leq i-p<r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n+1}}$
with $r_{t_{0}}:=0$. Since
$\alpha^{r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}}}(y)\in Y_{t_{n+1}},$
the identity
$\alpha^{i-p-(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}})}(\alpha^{r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}}}(y))=\alpha^{i-p}(y)\in Y$
implies that, since $0\leq i-p-(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}})<r_{t_{n+1}}$,
$i-p=r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}}\ \text{ and }\ q=t_{n+1}.$
Thus we have
$i=(r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{n}})+p=R_{y,k,n}+p$
and, by the assumption,
$\varsigma_{k}^{(i)}\circ\alpha^{-i}(x)$ $=\varsigma_{k}^{(i)}(y)$
$=\varsigma_{t_{n+1}}^{(p)}(\alpha^{R_{y,k,n}}(y))$
$=\varsigma_{t_{n+1}}^{(p)}(\alpha^{i-p}(y))$
$=\varsigma_{q}^{(p)}(\alpha^{i-p}(y))$
$=\varsigma_{q}^{(p)}\circ\alpha^{-p}(x).$
Similarly
$\varsigma_{l}^{(j)}\circ\alpha^{-j}(x)=\varsigma_{q}^{(p)}\circ\alpha^{-p}(x)$
can be obtained, and we see that $G$ is well defined.
Finally, if $x\in\alpha^{-j}(Y_{q})$ for $j=1,\ldots,m$, then $x\in\alpha^{i}(Y_{k})\cap\alpha^{-j}(Y_{q})$ for some $0\leq i<r_{k}$, $1\leq k\leq K$, hence $\alpha^{j}(x)\in\alpha^{i+j}(Y_{k})\cap Y_{q}\neq\emptyset$. Then $i+j\geq r_{k}$, and $i\geq r_{k}-j\geq r_{k}-m$. Therefore
$G(x)=\varsigma_{k}^{(i)}\circ\alpha^{-i}(x)=0,$
since $\varsigma_{k}^{(i)}=0$ for $i\geq r_{k}-m$. ∎
Corollary 8.4.
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. Then $\varsigma\in\pi(C(X))$ if and only if $\varsigma$ has the boundary decomposition property, and $\varsigma_{k}(x)\in\Gamma(\mathscr{M}_{k}|_{x})$ has a diagonal matrix representation for all $x\in\overline{Y_{k}}$ and $k=1,\dots,K$.
Proof.
Assume that $\varsigma$ has the boundary decomposition property and every $\varsigma_{k}(x)$ has a diagonal matrix representation. Then by Lemma 8.3 with $m=0$, there is a function $g\in C(X)=\Gamma(\mathscr{V}^{(0)})$ given by
$g\big{|}_{\alpha^{i}(\overline{Y_{k}})}=\varsigma_{k}^{(i)}\circ\alpha^{-i}\big{|}_{\alpha^{i}(\overline{Y_{k}})},\quad i=0,1,\dots,r_{k}-1.$
Thus $\varsigma_{k}^{(i)}(x)=g(\alpha^{i}(x))$ for $x\in\overline{Y_{k}}$, which means that $\varsigma_{k}=\pi_{k}(g)$ for all $k$. ∎
Corollary 8.5.
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ and $m$ be a positive integer. Then $\varsigma\in\pi(\mathcal{E}_{Y}^{\otimes m})$ if and only if $\varsigma$ satisfies the boundary decomposition property and has an $m$th lower subdiagonal matrix representation.
Proof.
Assume that $\varsigma$ satisfies the boundary decomposition property and has an $m$th (lower) subdiagonal matrix representation. Then $\varsigma_{k}(x)$ maps $\mathscr{V}_{x}^{(i)}$ in $\mathscr{D}^{(r_{k})}_{x}=\mathscr{V}_{x}^{(0)}\oplus\mathscr{V}_{x}^{(1)}\oplus\mathscr{V}_{x}^{(2)}\oplus\cdots\oplus\mathscr{V}_{x}^{(r_{k}-1)}$ to $\mathscr{V}_{x}^{(m+i)}$ for $i<r_{k}-m$ and $\{0\}$ for $r_{k}-m\leq i<r_{k}$. Again by Lemma 8.3, there is a $G\in\Gamma(\mathscr{V}^{(m)},\alpha^{m})$ such that
$G\big{|}_{\alpha^{i}(\overline{Y_{k}})}=\varsigma_{k}^{(i)}\circ\alpha^{-i}\big{|}_{\alpha^{i}(\overline{Y_{k}})},\quad i=0,1,2,\ldots,r_{k}-m-1,$
and vanishes on $\alpha^{-m}(Y)\cup\alpha^{-m+1}(Y)\cup\cdots\cup\alpha^{-1}(Y)$. By Proposition 3.2, there is $\xi\in\mathcal{E}_{Y}^{\otimes m}$ with $\psi(\xi)=G$. From the definition of $G$ we see that $\varsigma=\pi(\xi)$. ∎
Since $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ is generated (as a $C^{*}$-algebra) by elements that have $m$th (lower) subdiagonal matrix representation, we obtain the following theorem.
Theorem 8.6.
For $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$, $\varsigma$ satisfies the boundary decomposition property if and only if $\varsigma\in\pi(\mathcal{O}(\mathcal{E}_{Y}))$.
8.0.1. RSH structure of $\mathcal{O}(\mathcal{E}_{Y})$
Let $x\in\overline{Y_{2}}\setminus Y_{2}=\overline{Y_{2}}\cap Y_{1}$ be a point as in Definition 8.1. Then $r_{t_{s}}=r_{1}$ for all $s=1,\dots,m:=m(x,2)$ and thus $R_{x,2,s}=r_{t_{1}}+\cdots+r_{t_{s}}=sr_{1}$ for every $s$. Note that $m(x,2)=m$ for every $x\in\overline{Y_{2}}\setminus Y_{2}$.
Let $B_{1}:=\Gamma(\mathscr{M}_{1})$ and let $\rho_{1}:\Gamma(\mathscr{M}_{2})\to\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})$ be the restriction map. Define
$\varphi_{1}:\Gamma(\mathscr{M}_{1})\to\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})$
by $\varphi_{1}(\varsigma)(x):=\pi_{2}(\xi)(x)$ for $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ with $\pi_{1}(\xi)=\varsigma$. First the existence of such an element $\xi$ follows from the surjectivity of $\pi_{1}$ by Lemma 7.7. In order to see that $\varphi_{1}$ is well defined, we have to show that if $\pi_{1}(\xi_{1})=\pi_{1}(\xi_{2})=\varsigma$, then $\pi_{2}(\xi_{1})(x)=\pi_{2}(\xi_{2})(x)$ for all $x\in\overline{Y_{2}}\setminus Y_{2}$. Since $\varsigma=(\varsigma_{1},\dots,\varsigma_{K}):=\pi(\xi_{1})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ satisfies the boundary decomposition property by Theorem 8.6, we have from (23) that
$\pi_{2}(\xi_{1})(x)|_{\mathscr{D}_{x}^{(R_{x,2,s-1},r_{1})}}$ $\ =\varsigma_{2}(x)|_{\mathscr{D}_{x}^{(R_{x,2,s-1},r_{1})}}$
$\ =\varsigma_{1}(\alpha^{R_{x,2,s-1}}(x))\otimes id_{\mathscr{V}_{x}^{(R_{x,2,s-1})}}$
$\ =\pi_{1}(\xi_{1})(\alpha^{R_{x,2,s-1}}(x))\otimes id_{\mathscr{V}_{x}^{(R_{x,2,s-1})}}$
$\ =\pi_{1}(\xi_{2})(\alpha^{R_{x,2,s-1}}(x))\otimes id_{\mathscr{V}_{x}^{(R_{x,2,s-1})}}$
$\ =\pi_{2}(\xi_{2})(x)|_{\mathscr{D}_{x}^{(R_{x,2,s-1},r_{1})}}.$
Therefore, $\varphi_{1}$ is well defined, and obviously it is a -homomorphism. Let
$B_{2}:=$ $\ \Gamma(\mathscr{M}_{1})\oplus_{\Gamma(\mathscr{M}_{2}|_{\overline{Y_{2}}\setminus Y_{2}})}\Gamma(\mathscr{M}_{2})$
$=$ $\ \{(\varsigma_{1},\varsigma_{2})\in\Gamma(\mathscr{M}_{1})\oplus\Gamma(\mathscr{M}_{2})\mid\varphi_{1}(\varsigma_{1})=\rho_{1}(\varsigma_{2})\}.$
Then $B_{2}$ is an RSH algebra.
As in [16, Theorem 11.3.19], one can continue this process to form an RSH subalgebra $B_{k}$ of $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{k})$ for all $k=3,\dots,K$ simultaneously showing that at each step $k$, every element in $B_{k}$ satisfies the boundary decomposition property (only) up to $k$.
Theorem 8.7.
For $k=1,2,\ldots,K$, there is a recursive subhomogeneous algebra $B_{k}\subset\oplus_{i=1}^{k}\Gamma(\mathscr{M}_{i})$ such that $\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))=B_{k}$.
Proof.
We prove the theorem by induction on $k$. First note that by [32, Proposition 1.7] , $B_{1}:=\Gamma(\mathscr{M}_{1})$ is a recursive subhomogeneous algebra and $\pi_{1}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})$ is surjective by Lemma 7.7.
Assume that for $12\leq k\leq K-$ 1, the -homomorphism
$\oplus_{i=1}^{k-1}\pi_{i}:\mathcal{O}(\mathcal{E}_{Y})\to B_{k-1}\subset\oplus_{i=1}^{k-1}\Gamma(\mathscr{M}_{i})$
is surjective onto a recursive subhomogeneous algebra $B_{k-1}$. To define $B_{k}$ we need a -homomorphism
$\varphi_{k-1}:B_{k-1}\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}}).$
For $(b_{1},b_{2},\ldots,b_{k-1})\in B_{k-1}$, since $\oplus_{i=1}^{k-1}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))=B_{k-1}$, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ with $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},b_{2},\ldots,b_{k-1})$. Define
$\varphi_{k-1}(b_{1},b_{2},\ldots,b_{k-1}):=\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}.$
By Theorem 8.6, $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}$ is determined by $\pi_{i}(\xi)=b_{i}$, $1\leq i<k$, namely if $\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},\dots,b_{k-1})=\oplus_{i=1}^{k-1}\pi_{i}(\xi^{\prime})$ for $\xi$, $\xi^{\prime}\in\mathcal{O}(\mathcal{E}_{Y})$, then $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\pi_{k}(\xi^{\prime})|_{\overline{Y_{k}}\setminus Y_{k}}$. Hence $\varphi_{k-1}:B_{k-1}\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$ is a well-defined -homomorphism. Define $B_{k}$ as
$B_{k}:=B_{k-1}\oplus_{\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})}\Gamma(\mathscr{M}_{k})\subset\oplus_{i=1}^{k}\Gamma(\mathscr{M}_{i}),$
where $\rho_{k-1}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})$ is the restriction. Then $\oplus_{i=1}^{k}\pi_{i}(\mathcal{O}(\mathcal{E}_{Y}))$ is a $C^{*}$-subalgebra of the recursive subhomogeneous algebra $B_{k}$.
Now we prove that $\oplus_{i=1}^{k}\pi_{i}:\mathcal{O}(\mathcal{E}_{Y})\to B_{k}$ is surjective. Let
$(b_{1},b_{2},\ldots,b_{k})\in B_{k}=B_{k-1}\oplus_{\Gamma(\mathscr{M}_{k}|_{\overline{Y_{k}}\setminus Y_{k}})}\Gamma(\mathscr{M}_{k}).$
By induction hypothesis, there is $\xi\in\mathcal{O}(\mathcal{E}_{Y})$ such that
$\oplus_{i=1}^{k-1}\pi_{i}(\xi)=(b_{1},b_{2},\ldots,b_{k-1}).$
Consider the element
$(b_{1},b_{2},\ldots,b_{k})-\oplus_{i=1}^{k}\pi_{i}(\xi)=(0,0,\ldots,0,b_{k}-\pi_{k}(\xi)).$
Then the last coordinate $b_{k}-\pi_{k}(\xi)$ is not necessarily 00 over the whole $\overline{Y_{k}}$ but equal to 00 on $\overline{Y_{k}}\setminus Y_{k}$ since $\pi_{k}(\xi)|_{\overline{Y_{k}}\setminus Y_{k}}=\varphi_{k-1}(b_{1},b_{2},\ldots,b_{k-1})=\rho_{k-1}(b_{k})$. It is thus enough to claim that for any $\varsigma\in\Gamma(\mathscr{M}_{k})$ with $\varsigma|_{\overline{Y_{k}}\setminus Y_{k}}=0$, there is $\eta\in\mathcal{O}(\mathcal{E}_{Y})$ with
$\pi_{i}(\eta)=0\ \text{for }i=1,2,\ldots,k-1,\ \text{ and }\pi_{k}(\eta)=\varsigma.$
In fact, the claim implies that for $\varsigma:=b_{k}-\pi_{k}(\xi)\in\Gamma(\mathscr{M}_{k})$ there is such an $\eta\in\mathcal{O}(\mathcal{E}_{Y})$ and then $(b_{1},b_{2},\ldots,b_{k})=\oplus_{i=1}^{k}\pi_{i}(\xi+\eta)$ follows.
To prove the claim, let $\varsigma\in\Gamma(\mathscr{M}_{k})$ vanishes on $\overline{Y_{k}}\setminus Y_{k}$. We may assume that $\varsigma$ is $m$th lower subdiagonal in matrix representation, that is,
[…]
where $U\in\mathcal{U}$ is any open set with $x\in U$. Then for $i=0,\dots,r_{k}-1-m$, there exists $\varsigma^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that the map
$\varsigma(x)|_{\mathscr{V}^{(i)}_{x}}:\mathscr{V}^{(i)}_{x}\to\mathscr{V}^{(m+i)}_{x}=\mathscr{V}_{\alpha^{i}(x)}^{(m)}\otimes\mathscr{V}^{(i)}_{x}$
is given by $v\mapsto\varsigma^{(i)}(x)\otimes v$ for $v\in\mathscr{V}^{(i)}_{x}$. On $\overline{Y_{k}}\setminus Y_{k}$, all $\varsigma^{(i)}$’s are 0 by assumption of $\varsigma$. For notational convenience, set
$Y_{[i,j]}:=\overline{Y_{i}\cup\cdots\cup Y_{j}},\ \ i\leq j.$
Note that there is $j_{m}\in\{1,\dots,k\}$ for which
$r_{j_{m}}-1<r_{k}-1-m\leq r_{j_{m}+1}-1.$
In particular, if $m=0$, then $r_{j_{m}+1}=r_{k}$ (or $j_{m}=k-1$). Consider the following closed subset
$\mathcal{C}=$ $\,\cup_{i=0}^{r_{1}-1}\alpha^{i}(Y_{[1,k]})\,\cup\,\cup_{i=r_{1}}^{r_{2}-1}\alpha^{i}(Y_{[2,k]})\,\cup\,\cdots\,\cup\,\cup_{i=r_{j_{m}}}^{r_{k}-1-m}\alpha^{i}(Y_{[j_{m}+1,k]})$
$\,\cup\,\cup_{i=1}^{m}\alpha^{-i}(Y)$
of $X$. We will define a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ such that $\eta:=\psi^{-1}(\widetilde{\xi})$ satisfies the claim, where $\psi$ is the isomorphism of Proposition 3.2. For this, we first define a local section $\xi:\mathcal{C}\to\mathscr{V}^{(m)}|_{\mathcal{C}}$ and then extend it to $\widetilde{\xi}$ over the whole space $X$. Since the closed subsets $\alpha^{i}(Y_{[l,k]})$, $11\leq l\leq j_{m}+$ 1, $r_{l-1}\leq i\leq r_{l}-1$, of $\mathcal{C}$ have overlaps (and contain $\alpha^{i}(\overline{Y_{k}})$), we need $\xi$ to be well-defined on any possible overlaps.
Define, for $x\in\mathcal{C}$,
$\xi(x):=\begin{cases}\varsigma^{(i)}\circ\alpha^{-i}(x),&\text{for $x\in\alpha^{i}(\overline{Y_{k}})$},\\ 0,&\text{otherwise.}\end{cases}$
In order to check that $\xi(x)$ is well-defined on any possible overlaps, first let
$x=\alpha^{i}(z_{1})=\alpha^{j}(z_{2})\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{k}})$
for some $z_{1},z_{2}\in\overline{Y_{k}}$ and $i<j$. Then $\alpha^{j-i}(z_{2})=z_{1}\in Y$ while $0<j-i<r_{k}$. Thus $z_{2}\notin Y_{k}$, and so $z_{2}\in\overline{Y_{k}}\setminus Y_{k}$. Then $\varsigma^{(j)}\circ\alpha^{-j}(x)=\varsigma^{(j)}(z_{2})=0$. Also $\alpha^{r_{k}-j+i}(z_{1})=\alpha^{r_{k}}(z_{2})\in Y$ implies $z_{1}\notin Y_{k}$ since $0<r_{k}-j+i<r_{k}$. Then $z_{1}\in\overline{Y_{k}}\setminus Y_{k}$ and similarly we have $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(z_{1})=0$. Thus $\xi$ is well-defined (to be 00) on overlaps of the form $\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{k}})$. To show that $\xi$ is well defined to be 00 on an overlap
$\alpha^{i}(\overline{Y_{k}})\cap\big{(}\cup_{j=1}^{m}\alpha^{-j}(Y)\big{)}$
for $0\leq i\leq r_{k}-1-m$, suppose that $x=\alpha^{i}(\overline{y_{k}})\in\cup_{j=1}^{m}\alpha^{-j}(Y)$ for some $\overline{y_{k}}\in\overline{Y_{k}}$. Then $\alpha^{j+i}(\overline{y_{k}})\in Y$ for some $j$, but $10\leq j+i\leq r_{k}-$ 1 hence $\overline{y_{k}}\notin Y_{k}$. Then $\overline{y_{k}}\in\overline{Y_{k}}\setminus Y_{k}$ again, and then $\varsigma^{(i)}\circ\alpha^{-i}(x)=\varsigma^{(i)}(\overline{y_{k}})=0$. Now suppose $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(Y_{k})\neq\emptyset$ for some $1\leq q<k$, $0\leq p<r_{q}$, $0\leq i<r_{k}$, with an element $x=\alpha^{p}(\overline{y_{q}})=\alpha^{i}(y_{k})$ where $\overline{y_{q}}\in\overline{Y_{q}}$ and $y_{k}\in Y_{k}$. Then $p\geq i$ by (26), and $\alpha^{r_{q}-p+i}(y_{k})=\alpha^{r_{q}-p}(x)=\alpha^{r_{q}-p}(\alpha^{p}(\overline{y_{q}}))=\alpha^{r_{q}}(\overline{y_{q}})\in Y$, which is not possible because $0<r_{q}-p+i<r_{k}$ and $y_{k}\in Y_{k}$. Thus $\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}})=\alpha^{p}(\overline{Y_{q}})\cap\alpha^{i}(\overline{Y_{k}}\setminus Y_{k})$ on which $\varsigma^{(i)}\circ\alpha^{-i}=0$. Thus $\xi$ is well-defined over the closed subset $\mathcal{C}$ of $X$. Using the Tietze extension theorem for vector bundles, we further extend the domain of $\xi$ to the whole $X$ and get a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over $X$ such that
$\widetilde{\xi}|_{\alpha^{i}(\overline{Y_{q}})}=\xi|_{\alpha^{i}(\overline{Y_{q}})}=0$
for $11\leq q\leq k-$ 1, $0\leq i\leq\min\{r_{q}-1,r_{k}-1-m\}$,
$\widetilde{\xi}|_{\alpha^{i}(\overline{Y_{k}})}=\xi|_{\alpha^{i}(\overline{Y_{k}})},$
for $0\leq i\leq r_{k}-1-m$, and $\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0$. Under the isomorphism $\psi:\mathcal{E}^{\otimes m}=\Gamma(\mathscr{V},\alpha)^{\otimes m}\to\Gamma(\mathscr{V}^{(m)},\alpha^{m})$, the property $\widetilde{\xi}|_{\cup_{j=1}^{m}\alpha^{-j}(Y)}=0$ implies that $\eta=\psi^{-1}(\widetilde{\xi})\in\mathcal{E}_{Y}^{\otimes m}=E_{m}\subset\mathcal{O}(\mathcal{E}_{Y})$. If $m=0$, then $\cup_{j=1}^{m}\alpha^{-j}(Y):=\emptyset$, $\mathcal{E}^{\otimes 0}=\Gamma(\mathscr{V}^{(0)},\alpha^{0})=C(X)$ and $\psi$ is the identity -homomorphism of $C(X)$.
Finally, for $x\in\overline{Y_{i}}$, $11\leq i\leq k-$ 1, and a vector $(a_{0},a_{1},\ldots,a_{r_{i}-1})\in\mathscr{D}^{(r_{i})}_{x}=\mathscr{V}^{(0)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{i}-1)}_{x}$, we have
$\ \pi_{i}(\psi^{-1}(\widetilde{\xi}))(x)(a_{0},a_{1},\ldots,a_{r_{i}-1})$
$=$ $\ (0,\ldots,0,\psi(\psi^{-1}(\widetilde{\xi}))(x)\otimes a_{0},\psi(\psi^{-1}(\widetilde{\xi}))(\alpha(x))\otimes a_{1},$
$\qquad\qquad\qquad\qquad\ldots,\psi(\psi^{-1}(\widetilde{\xi}))(\alpha^{r_{i}-1-m}(x))\otimes a_{r_{i}-1-m})$
$=$ $\ (0,\ldots,0,\widetilde{\xi}(x)\otimes a_{0},\widetilde{\xi}(\alpha(x))\otimes a_{1},\ldots,\widetilde{\xi}(\alpha^{r_{i}-1-m}(x))\otimes a_{r_{i}-1-m})$
$=$ $\ (0,\ldots,0,0,\ldots,0)=0,$
and for $x\in\overline{Y_{k}}$, $\pi_{k}(\psi^{-1}(\widetilde{\xi}))=\varsigma$ (see Remark 7.5). Hence $\pi_{i}(\eta)=0$ for $1\leq i\leq,k-1$ and $\pi_{k}(\eta)=\varsigma$ as desired. ∎
Remark 8.8.
  1. (1)
    Let $y\in X$ be a single point, and let $Y_{1}\supset Y_{2}\supset Y_{3}\supset\cdots$ be a sequence of closed subsets of $X$, each with non-empty interiors, satisfying
    $\cap_{n\in\mathbb{Z}_{>0}}Y_{n}=\{y\}.$
    Then
    $\mathcal{O}(\mathcal{E}_{\{y\}})\cong\underset{\longrightarrow}{\lim}\mathcal{O}(\mathcal{E}_{Y_{n}})$
    and so, by the above, $\mathcal{O}(\mathcal{E}_{\{y\}})$ is an approximately subhomogeneous (ASH) $\mathrm{C}^{*}$-algebra. The topological dimension of the RSH decomposition of $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is bounded by the dimension of $X$. Thus if $X$ is finite dimensional, $\mathcal{O}(\mathcal{E}_{\{y\}})$ is an ASH with no dimension growth. It follows that $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is $\mathcal{Z}$-stable and thus classifiable. Since $\mathcal{O}(\mathcal{E}_{Y_{n}})$ is a centrally large subalgebra of $\mathcal{O}(\mathcal{E})$, it too is $\mathcal{Z}$-stable and hence classifiable. Thus we recover Theorem 6.18 of [2].
  2. (2)
    If $(X,\alpha)$ is only assumed to have mean dimension zero (which means $\dim X$ may be infinite), the techniques above do not apply. However, using the RSH decomposition of $\mathrm{C}^{*}(C(X),C_{0}(X\setminus Y)u)\subset C(X)\rtimes_{\alpha}X$ for $Y\subset X$ a closed subset with non-empty interior, Elliott and Niu show that for any point $y\in X$ $\mathrm{C}^{*}(C(X),C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}X$ can be locally approximated by subhomogeneous $\mathrm{C}^{*}$-algebras with arbitrarily slow dimension ratio [15], which is enough to imply $\mathcal{Z}$-stability (and hence classifiability) of both $\mathrm{C}^{*}(C(X),C_{0}(X\setminus\{y\})u)$ and $C(X)\rtimes_{\alpha}X$.
  3. (3)
    Let $(X,\alpha)$ be a minimal dynamical system. In [3], it is shown that crossed product $C(X)\rtimes_{\alpha}\mathbb{Z}$ has stable rank one, with no restrictions an $(X,\alpha)$. In particular, $C(X)\rtimes_{\alpha}\mathbb{Z}$ may have stable rank one even if it is not $\mathcal{Z}$-stable. This is shown by exhibiting a so-called diagonal subhomogeneous (DSH) algebra decomposition for orbit-breaking subalgebras when $Y\subset X$ is closed with non-empty interior. A DSH algebra is a special type of RSH algebra where the gluing maps between stages of the RSH decomposition are assumed to be diagonal (see [3, Definition 2.3] for a precise definition). As above, if $y$ is a single point, then this implies $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)\subset C(X)\rtimes_{\alpha}\mathbb{Z}$ is approximately DSH. They show that a simple inductive limit of DSH algebras with diagonal connecting maps has stable rank one [3, Theorem 3.30]. Thus $\mathrm{C}^{*}(X,C_{0}(X\setminus\{y\})u)$ has stable rank one, and as it is a centrally large subalgebra, this passes to $C(X)\rtimes_{\alpha}\mathbb{Z}$. The RSH constructions in this paper will not be DSH algebra in general. This follows from [3, Corollary 2.22]. Nevertheless, our RSH constructions should be amenable to similar techniques. The question of stable rank of $\mathrm{C}^{*}$-algebras associated to homeomorphisms twisted by (non-trivial) line bundles is the subject of future investigations.